### v2

# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (BRIGHT) to support AI-based all-weather disaster response. To the best of our knowledge, BRIGHT is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of man-made disasters across 14 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in BRIGHT, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In the experiments, we have tested seven advanced AI models trained with BRIGHT to validate the transferability and robustness. Beyond that, BRIGHT also serves as a challenging benchmark for a variety of tasks under disaster scenarios, including unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM). The experimental results serve as baselines to inspire future research and model development. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as new disaster data becomes available. BRIGHT also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven EO methods in support of people in disaster-affected areas.
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods have affected 2.85 billion people and caused approximately 1.12 million deaths. These disasters have also resulted in economic losses of US$2,647 billion in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions to minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, preventing efficient on-site assessments. Earth observation (EO), also known as remote sensing, provides a safe and efficient way to obtain information on building damage in disaster areas because of its wide field of view and contactless operation.
The EO technologies commonly used for assessing building damage after disasters are optical and synthetic aperture radar (SAR). Optical imagery is a primary source for building damage assessment because of its intuitive and easy-to-interpret nature. For example, moderate-resolution optical EO data from the Landsat series and Sentinel-2 have been used to assess building damage (Yusuf et al., 2001; Fan et al., 2019; Sandhini Putri et al., 2022). Landsat and Sentinel-2 data are limited in spatial resolution and only provide broad approximations of affected areas, which lack accuracy for specific buildings crucial for timely rescue. The new generation of very-high-resolution (VHR) optical sensors, such as IKONOS and WorldView, provide EO data with spatial resolutions at a meter and a sub-meter level, enabling finer assessments at the level of individual buildings (Freire et al., 2014). These data have been used successfully for the assessment of building damage (Yamazaki and Matsuoka, 2007; Tong et al., 2012; Freire et al., 2014).
While accurate building damage maps can be obtained by human-expert visual interpretation of optical images, this process is time-consuming and labor-intensive for rapid large-scale assessments. In addition, it requires trained professionals. Therefore, recent literature has focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, deep learning (DL) techniques have significantly improved the efficiency and accuracy of building damage assessment. Earlier work was toward a single disaster event with labels annotated for a particular disaster area to train a model, which is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) of different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Shen et al., 2022; Guo et al., 2024; Wang et al., 2024; Chen et al., 2022a; Kaur et al., 2023; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damages of buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1c is not affected by smoke and clearly shows the buildings damaged by the wildfire. Due to the advantages of SAR imagery, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate accurate building damage mapping compared to optical imagery (Brunner et al., 2010; Adriano et al., 2021). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This is a significant challenge that remains and needs to be addressed.
To support DL research aimed at all-weather building damage mapping, we present BRIGHT, the first open and globally distributed multimodal VHR dataset for building damage assessment. Advances in EO technology have enabled data providers like Capella111https://www.capellaspace.com/ and Umbra222https://umbra.space/ to offer VHR SAR imagery at a sub-meter level resolution per pixel. This allows for detailed building damage assessments at the individual building level to guide targeted and effective rescue operations as required by emergency responders. Benefiting from the progress that has been made in EO, BRIGHT incorporates both pre-event optical imagery and post-event SAR imagery with spatial resolutions ranging from 0.3 meters to 1 meter per pixel. The types of disaster events considered in BRIGHT are earthquakes, storms (e.g., hurricane, cyclone), wildfires, floods, and volcanic eruptions. These natural disasters accounted for 84 of the fatalities and 94 of the economic losses between 1998 and 2017 (UNDRR, 2018a). In addition to natural disasters, the BRIGHT dataset further considers disasters caused by human activity, such as accidental explosions and armed conflicts, which also pose significant threats to human life and infrastructure and can occur unexpectedly, requiring a rapid response (UNDRR, 2018b). The events cover 14 different regions distributed around the globe with a focus on developing countries where external assistance is most urgently needed after a disaster. The labels are manually annotated with multi-level annotations that distinguish between damaged buildings and completely destroyed buildings.
| Dataset | OA | Modality | GSD (m/pixel) | No. of events | Disaster type | No. of building | Granularity |
| ABCD (Fujita et al., 2017) | Yes | Optical EO | 0.4 | 1 | Tsunami | N/A | Image-level |
| (Nguyen et al., 2017) | Yes | Images on social media | N/A | 4 | 3 natural disasters | N/A | Image-level |
| (Cheng et al., 2021) | Yes | UAV image | N/A | 1 | Hurricane | 1,802 | Image-level |
| (Xue et al., 2024) | Yes | Street-view image | N/A | 1 | Hurricane | 2,468 | Image-level |
| QQB (Sun et al., 2024b) | Yes | Optical and SAR EO | <1 | 1 | Earthquake | 4,029 | Pixel-level |
| FloodNet (Rahnemoonfar et al., 2021) | Yes | UAV image | N/A | 1 | Flood | 6,675 | Pixel-level |
| RescueNet (Rahnemoonfar et al., 2023) | Yes | UAV image | N/A | 1 | Hurricane | 10,903 | Pixel-level |
| Ida-BD (Kaur et al., 2023) | No | Optical EO | 0.5 | 1 | Hurricane | 18,083 | Pixel-level |
| xBD (Gupta et al., 2019) | Yes | Optical EO | <0.8 | 15 | 6 natural disasters | >700,000 | Pixel-level |
| BDD (Adriano et al., 2021) | No | Optical and SAR EO | 1.2-3.3 | 9 | 3 natural disasters | 123,453 | Pixel-level |
| BRIGHT | Yes | Optical and SAR EO | 0.3-1 | 14 | 5 natural disasters 2 human-made disasters | 384,596 | Pixel-level |
0.1 Comparison with existing datasets
The comparison between BRIGHT and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to limited available disaster events with corresponding open-source EO data and annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. The xBD includes only optical EO data. Therefore, it does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models. The dataset most similar to BRIGHT is the BDD proposed by Adriano et al. (2021). The main differences between BDD and BRIGHT datasets are: (1) BRIGHT covers more disaster events and building instances, including both natural and human-made disasters. (2) BRIGHT has higher spatial resolution; while the highest resolution of SAR images in BDD is 1.2 meters, BRIGHT provides finer detail with spatial resolution from 0.3 meters to 1 meter to enable the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, BRIGHT is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not concerned with building damage assessment.
0.2 Main contribution
The contributions of this paper are twofold:
1.
We present BRIGHT, the first multimodal building damage dataset with a spatial resolution greater than 1 meter, which is publicly available to the community. BRIGHT employs a combination of pre-event optical imagery and post-event SAR imagery with various disaster events and rich geographic diversity to support the study of AI-based multimodal building damage mapping, especially in developing countries.
2.
We evaluate a suite of contemporary models on BRIGHT to establish robust baselines. Beyond supervised deep learning, BRIGHT can support a wide range of AI-based methods. It enables research in UDA, SSL, UMCD, and UMIM, among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with source code and pretrained model weights, are publicly released to provide strong baselines and accelerate future developments in EO-based disaster response.
1 Study areas and disaster events
We selected 14 disaster events across the globe for our BRIGHT dataset, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts.
| Disaster area | Type of disaster | Date | GSD (m/pixel) | Provider | No. of tiles | No. of building |
| Beirut, Lebanon | Explosion (EP) | 04 Aug. 2020 | 1 | Maxar Capella | 133 | 25,496 |
| Bata, Equatorial Guinea | Explosion (EP) | 07 Mar. 2021 | 0.5 | Maxar Capella | 107 | 8,893 |
| Goma, DR Congo | Volcano eruption (VE) | 22 May 2021 | 0.33 | Maxar Capella | 123 | 18,741 |
| Les Cayes, Haiti | Earthquake (EQ) | 14 Aug. 2021 | 0.48 | Maxar Capella | 73 | 18,918 |
| La Palma, Spain | Volcano Eruption (VE) | 19 Sept. 2021 - 13 Dec. 2021 | 0.3-0.35 | IGN (Spain) Capella | 933 | 30,239 |
| Boulder, USA | Wildfire (WF) | 30 Dec. 2021 - 01 Jan., 2022 | 0.6 | NAIP Capella | 77 | 8,365 |
| Ukraine | Armed conflict (AC) | 22 Mar. 2022 - 21 Sept. 2022 | 0.6 | Google Earth Capella | 513 | 56,770 |
| Turkey | Earthquake (EQ) | 06 Feb. 2023 | 0.30-0.35 | Maxar Capella Umbra | 1,114 | 135,033 |
| Kyaukpyu, Myanmar | Cyclone (CC) | 14 May 2023 | 0.6 | Google Earth Capella | 126 | 8,052 |
| Maui, Hawaii, USA | Wildfire (WF) | 08 Aug. 2023 - 09 Aug. 2023 | 0.6 | NOAA Capella | 65 | 3,995 |
| Morocco | Earthquake (EQ) | 08 Sept. 2023 | 0.35-0.4 | Maxar Capella | 567 | 6,269 |
| Derna, Libya | Flood (FL) | 10 Sept. 2023 | 0.35 | Maxar Capella | 124 | 1,0979 |
| Acapulco, Mexico | Hurricane (HC) | 25 Oct. 2023 | 0.35-0.8 | Google Earth Capella | 212 | 18,437 |
| Noto, Japan | Earthquake (EQ) | 01 Jan. 2024 | 0.5 | GSI Umbra | 79 | 8,153 |
| Total | - | - | 0.3-1 | - | 4,246 | 384,596 |
1.1 Explosion in Beirut, 2020
On August 4, 2020, a massive explosion occurred at the Port of Beirut in Lebanon, caused by the improper storage of 2,750 tons of ammonium nitrate. The explosion caused widespread damage within a radius of several kilometers, significantly impacting the port and the surrounding neighborhoods, including areas such as Gemmayzeh, Mar Mikhael, and Achrafieh. It resulted in 218 deaths, more than 7,000 injuries, and left approximately 300,000 people homeless. Economic losses were estimated to be between 10 billion and 15 billion USD. The disaster compounded Lebanon’s ongoing economic challenges and contributed to political instability and social unrest.
1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city. A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
1.3 Volcano Eruption in DR Congo and Rwanda, 2021
On May 22, 2021, Mount Nyiragongo in the Democratic Republic of the Congo erupted, causing widespread devastation. The eruption resulted in 32 deaths, the destruction of 1,000 homes, and the displacement of thousands as lava flows threatened the city of Goma. Nearly 400,000 people were evacuated due to the risk of further volcanic activity, including potential magma flow beneath Goma and nearby Lake Kivu. Despite continued seismic activity, life in Goma largely returned to normal by August 2021, although plans to relocate parts of the city remain under consideration due to the ongoing threat from the volcano.
1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed. As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
1.5 Volcano Eruption in La Palma, 2021
On September 19, 2021, the Cumbre Vieja volcano on La Palma, part of Spain’s Canary Islands, erupted following several days of seismic activity. The eruption primarily impacted the island’s western side, covering over 1,000 hectares with lava and destroying more than 3,000 buildings, including the towns of Todoque and La Laguna. The lava flow, measuring about 3.5 kilometers wide and 6.2 kilometers long, reached the sea, cutting off the coastal highway and forming a new peninsula with extensive lava tubes. Although the timely evacuation of around 7,000 people prevented major casualties, one person died from inhaling toxic gases. Economic losses exceeded 800 million Euros, and thousands of residents were displaced, with significant damage to arable land and livelihoods.
1.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the state history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over US$2 billion in damage. More than 37,500 residents evacuated, and significant damage to public drinking water systems occurred.
1.7 Armed Conflict in Ukraine, 2022
In February 2022, the conflict in Ukraine escalated significantly when Russian forces launched a full-scale invasion. The invasion has affected many parts of Ukraine, especially the eastern regions and the major cities, causing widespread destruction to homes, infrastructure, and industrial facilities. The conflict has resulted in thousands of civilian and military casualties and has caused the displacement of millions of people. Out of a population of 41 million, about 8 million Ukrainians were internally displaced, while more than 8.2 million had fled the country by April 2023, creating Europe’s largest refugee crisis since World War II. The economic losses are substantial, with extensive damage to infrastructure, industrial output, and the broader economy.
1.8 Earthquake in Turkey, 2023
On February 6, 2023, a magnitude 7.8 earthquake struck southeastern Turkey near Gaziantep, followed by a magnitude 7.7 aftershock. The disaster, the most powerful earthquake in Turkey since 1939, caused widespread destruction across approximately 350,000 , affecting 14 million people and displacing 1.5 million. The death toll reached 53,537 in Turkey and 5,951 in Syria, with 107,213 injuries, making it one of the deadliest earthquakes in modern history. Economic losses were estimated at USD 148.8 billion in Turkey and USD 14.8 billion in Syria, with over 518,000 houses and 345,000 apartments destroyed. The earthquake caused severe damage to infrastructure, agriculture, and essential services, further worsening the region’s economic challenges. International aid was mobilized to support the affected populations.
1.9 Cyclone in Myanmar, 2023
In May 2023, Cyclone Mocha, a Category 4 hurricane, struck Myanmar, causing widespread devastation in the country’s coastal regions, particularly in Rakhine State. The storm resulted in at least 460 deaths, with over 700 people injured, mostly among the Rohingya refugees. The cyclone destroyed over 183,000 houses, 1,770 religious buildings, and 1,397 schools, and caused US$2.24 billion in damages, equivalent to 3.4% of Myanmar’s GDP. Severe flooding affected 895 of land, displacing over 1.2 million people. In Rakhine, towns like Sittwe were heavily flooded, with a storm surge of up to 3.5 meters, while nearly 90% of homes in Kyauktaw were destroyed. The cyclone severely affected communication networks and infrastructure, exacerbating the region’s vulnerability to future natural disasters.
1.10 Wildfire in Hawaii, 2023
In August 2023, a series of wildfires broke out on the island of Maui, Hawaii, causing widespread destruction and significant impacts on the local population and environment. The fires, fueled by dry conditions and strong winds, primarily affected the town of Lahaina, where at least 102 people were killed and two remain missing. Over 2,200 buildings were destroyed, including many historic landmarks, resulting in estimated damages of US$5.5 billion. The fires prompted evacuations and led to the displacement of thousands of residents, with significant economic losses in the tourism and agriculture sectors.
1.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The quake resulted in at least 2,960 deaths and 5,674 injuries, affecting over 2.8 million people, including 100,000 children. The earthquake damaged or destroyed 40,759 houses and 2,930 villages, with 19,095 additional houses collapsing. Significant destruction occurred in rural areas of the Atlas Mountains. Economic losses could reach up to 9% of Morocco’s GDP. Historic sites in Marrakesh, including parts of the Medina and several mosques, were severely damaged.
1.12 Flood and Storm in Libya, 2023
In September 2023, northeastern Libya was hit by Storm Daniel, causing catastrophic flooding, particularly in the city of Derna. The storm led to the collapse of two dams, releasing 30 million cubic meters of water and partially destroying the city. Casualty estimates range from 5,900 to 20,000, making it the second deadliest dam failure in history. The flooding affected seven out of 10 districts in Derna and left more than 40,000 people displaced. More than 2,200 buildings were flooded and the collapse of four bridges split the city into two. The economic losses were substantial, with severe damage to infrastructure, homes, and the agricultural sector.
1.13 Hurricane in Mexico, 2023
In September 2023, Hurricane Norma, a Category 4 hurricane, struck the western coast of Mexico, severely affecting Sinaloa and Baja California Sur. The storm caused over 100 deaths and displaced thousands, with significant damage to homes, infrastructure, and agriculture. This was followed by Hurricane Otis in October, which made landfall near Acapulco as a Category 5 hurricane. Otis was the strongest Pacific hurricane to hit Mexico, causing at least 52 deaths and leaving 32 missing. The storm caused unprecedented destruction, with more than 51,864 homes destroyed and damages estimated at US$12–16 billion, surpassing Hurricane Wilma as the costliest Mexican hurricane.
1.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 6.58-meter tsunami along the Sea of Japan coast. The disaster resulted in 339 deaths and over 1,300 injuries, with 156,526 structures damaged across nine prefectures. It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20, 12,929 people remained in 521 evacuation centers. The event prompted Japan’s first major tsunami warning since 2011. Economic losses are estimated to be substantial, with significant impacts on infrastructure and communities across the affected regions.
2 Dataset Description
2.1 Construction of the dataset
Figure 3 shows the flowchart of developing BRIGHT. The optical EO data in our dataset are mainly from Maxar’s Open Data program333https://www.maxar.com/open-data, while the SAR EO data are from Capella Space444https://www.capellaspace.com/earth-observation/gallery and Umbra Space555https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consist of red, blue, and green bands, while the SAR EO data are amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the data to an 8-bit data format. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two types of EO data. Figure 11 shows the selected control points on three disaster scenes.
| Class | Definition |
| Background (0) | All non-building pixels |
| Intact (1) | No visible signs of structural damage, water intrusion, shingle displacement, or burn marks. |
| Damaged (2) | Partial structural damage to the building, such as missing roof members, visible cracks, or partial wall/roof collapse. Buildings may be partially burned, surrounded by water or mud, or affected by nearby volcanic flows. |
| Destroyed (3) | Completely collapsed, burned, partially/completely covered by water/mud or no longer present |
The labels in BRIGHT consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service666https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products777https://unosat.org/products, and The Federal Emergency Management Agency (FEMA)888https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To further improve damage information reliability, we performed a secondary validation and correction of these annotations. EO experts carefully reviewed the damage points using multi-temporal imagery available on platforms such as Google Earth and manually refined the labels when inconsistencies or misclassifications were observed. Since the number of damage levels varied across different events, we refered to the framework in FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019) to standardize them into three categories: intact (1), damaged (2), and destroyed (3). Table 3 summarizes these definitions. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to possible coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
Figure 4 presents thumbnails of selected local areas from the 14 disaster events included in the BRIGHT dataset.
2.2 Statistics for BRIGHT
The basic information about BRIGHT, including disaster events, EO data, the number of corresponding EO tiles, and the total number of building pixels, is summarized in Table 2. After cropping the EO data into 10241024-pixel tiles, BRIGHT contains 4,246 multimodal image pairs.
The key statistics of BRIGHT are illustrated in Figure 5. Figure 5a shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting BRIGHT’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained deep learning (DL) models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5b shows that BRIGHT covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
Figure 5c further illustrates the feature distribution of buildings in optical and SAR images for the same four events shown in Figure 5a, demonstrating clear inter-event separability in both modalities. Finally, BRIGHT faces a significant challenge of sample imbalance. There is a notable imbalance between background pixels and foreground (building) pixels, with a ratio of approximately 7:1. Additionally, an imbalance exists within damage categories: only about 6.5 of building pixels represent destroyed buildings, 6.7 correspond to damaged buildings, and 86.8 are intact buildings. This imbalance complicates model training, necessitating careful strategies to develop robust DL models.
| Descriptor | RMSE (All) | RMSE (Building) |
| RIFT (Li et al., 2020) | 1.125 | 1.138 |
| LNIFT (Li et al., 2022) | 0.857 | 0.825 |
| SRIF (Li et al., 2023a) | 1.090 | 1.054 |
| Average | 1.024 | 1.006 |
We also evaluated the registration accuracy between optical and SAR EO data in BRIGHT. Due to the absence of real ground truth, we introduce a proxy method to estimate registration errors, leveraging existing multimodal image descriptors. This approach is detailed in Appendix B. Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements, obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within building regions. Accurate registration results ensure spatial consistency across modalities, providing a solid foundation for accurate building damage assessment in BRIGHT.
2.3 Dataset division strategy
To train ML/DL models using BRIGHT and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
In addition to the above standard ML division, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events, a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. This setup comprises two key divisions:
Zero-shot setup: This setting mimics a real-world scenario where a newly occurring disaster must be analyzed without any prior labeled data from the same event. We isolate one event as an unseen test set while using the remaining 13 events for training and validation. This setting evaluates the cross-event generalization ability of models, testing how well they can transfer learned knowledge from previous disasters to an entirely new event. Due to the high variability of disaster types and geographies, this setup is inherently challenging, as models trained on past disasters may struggle to assess damage patterns in a previously unseen event accurately.
One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup, acknowledging that in real-world disaster response, a small number of labeled samples from the new disaster can be obtained. In this setting, a limited subset of labeled data (one image for training and one image for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.
3 Methodology
3.1 Problem statement and two solutions
The objective of building damage assessment is to interpret EO data covering the affected area and generate a building damage proxy map that reflects the extent of damage to buildings. To achieve this, two common solutions are typically employed.
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted, which can be formalized as , where is is the pre-event imagery, is the post-event imagery, is a semantic segmentation model, is the obtained damage map. In the context of this paper, is VHR optical imagery and is VHR SAR imagery.
The second solution adopts the task decoupling approach (Gupta et al., 2019; Zheng et al., 2021), which breaks down building damage assessment into two subtasks: the building localization task, i.e., separating the building from the background, and the damage classification task, i.e., focusing on the classification between different levels of damage. This solution can be formulated as and , where is the building localization map, is the damage classification map, and are models for building localization and damage classification tasks, respectively. and can be two separate models (Gupta et al., 2019) or a unified multi-task learning model (Zheng et al., 2021; Chen et al., 2022a, 2024). The final building damage map is obtained by combining the two outputs using a simple mask operation: .
Since the aim of this article is not only to provide a large-scale multimodal benchmark dataset to support all-weather disaster response but also to offer insights for designing appropriate methods in future research, we will employ and compare both solutions in our experiments.
3.2 Benchmark suites
We evaluate BRIGHT using several advanced deep network architectures from both the computer vision and remote sensing communities. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of representative deep learning paradigms, including CNNs, RNNs, Transformers, and the more recent Mamba architecture.
Among these seven networks, UNet, DeepLabV3+, and SiamAttnUNet adopt the first solution, directly treating building damage assessment as a single semantic segmentation task. In contrast, SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba adopt the second solution by decoupling the task into two subtasks: building localization and damage classification.
Beyond supervised deep learning, BRIGHT also enables the evaluation of broader strategies commonly explored in EO and vision communities:
Unsupervised domain adaptation (UDA) methods for the zero-shot transfer setup, enabling models to transfer knowledge across disaster events with no labeled samples from the target event.
Semi-supervised learning (SSL) approaches for the one-shot transfer setup, leveraging a small number of labeled samples and the remaining unlabeled samples from new disaster events to refine model adaptation.
Unsupervised multimodal change detection (UMCD) methods, which exploit the modality-independent relationship in optical and SAR data to detect land-cover changes without requiring manual annotations.
Unsupervised multimodal image matching (UMIM) methods, which aim to learn modality-independent features to enable automatic registration of multimodal data without relying on manual alignment.
| Method | () | () | Final OA () | Final mIoU () | IoU per class () | |||
| Background | Intact | Damaged | Destroyed | |||||
| UNet | 87.97 | 72.24 | 95.47 | 64.94 | 96.19 | 71.27 | 39.13 | 53.17 |
| DeepLabV3+ | 87.00 | 70.33 | 95.43 | 64.80 | 95.98 | 70.98 | 37.53 | 54.69 |
| SiamAttnUNet | 88.16 | 70.13 | 95.45 | 64.26 | 96.14 | 71.90 | 35.06 | 53.92 |
| SiamCRNN | 89.45 | 72.02 | 95.76 | 65.73 | 96.48 | 73.44 | 38.91 | 54.18 |
| ChangeOS | 89.60 | 71.88 | 95.84 | 65.98 | 96.54 | 73.85 | 38.99 | 54.53 |
| DamageFormer | 90.29 | 72.51 | 96.13 | 67.09 | 96.87 | 75.04 | 39.86 | 56.59 |
| ChangeMamba | 90.90 | 72.70 | 96.22 | 67.63 | 96.96 | 75.59 | 40.05 | 57.91 |
3.3 Model training
To train these supervised models, we use a combination of cross-entropy loss and Lovasz softmax loss (Berman et al., 2018). Cross-entropy loss serves as the basic loss function for dense prediction tasks, while Lovasz softmax loss effectively addresses sample imbalance between non-building and building pixels and across different damage levels.
For UNet, DeepLabV3+, and SiamAttnUNet, which directly predict damage proxy maps from the input multimodal image pairs, the training loss function is defined as:
For SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba, which decouple building damage assessment into building localization and damage classification subtasks, the training loss function is defined as:
All models are trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of 1 and a weight decay of 5. The training process consists of 50,000 iterations, with a batch size of 16. To enhance sample diversity and improve model generalization, we apply several data augmentation techniques, including random flipping, random rotation (in 90-degree increments), and random cropping.
For zero-shot and one-shot setups, we incorporate UDA and SSL techniques to better utilize limited or unlabeled target disaster data. The specifics of these training adjustments, along with the UMCD and UMIM approaches, are provided in Appendix E, F, and G, respectively.
| Events | UNet | DeepLabV3+ | SiamAttnUNet | SiamCRNN | ChangeOS | DamageFormer | ChangeMamba |
| Beirut-EP-2020 | 52.23 | 48.53 | 51.85 | 52.04 | 52.57 | 60.55 | 49.38 |
| Bata-EP-2021 | 39.24 | 40.26 | 38.69 | 41.44 | 39.80 | 40.37 | 40.47 |
| Goma-VE-2021 | 61.52 | 59.26 | 62.95 | 62.50 | 61.84 | 61.55 | 62.89 |
| Les Cayes-EQ-2021 | 40.39 | 38.52 | 40.84 | 40.99 | 41.24 | 41.75 | 42.36 |
| La Palma-VE-2021 | 62.72 | 64.31 | 65.56 | 64.24 | 65.34 | 64.27 | 65.13 |
| Marshall-WF-2021 | 53.96 | 44.67 | 51.48 | 56.57 | 56.55 | 57.50 | 57.23 |
| Ukraine-AC-2022 | 43.92 | 43.84 | 40.71 | 45.52 | 45.56 | 47.45 | 47.43 |
| Turkey-EQ-2023 | 49.34 | 49.53 | 48.57 | 50.69 | 52.34 | 51.36 | 52.70 |
| Kyaukpyu-CC-2023 | 38.11 | 38.56 | 40.49 | 42.35 | 42.03 | 48.42 | 41.05 |
| Hawaii-WF-2023 | 53.02 | 57.81 | 57.92 | 55.34 | 58.40 | 60.82 | 60.70 |
| Morocco-EQ-2023 | 42.62 | 42.05 | 42.58 | 43.17 | 43.05 | 44.00 | 44.06 |
| Derna-FL-2023 | 59.47 | 56.08 | 66.82 | 59.48 | 57.73 | 62.24 | 63.30 |
| Acapulco-HC-2023 | 49.21 | 46.66 | 42.15 | 49.98 | 51.08 | 50.82 | 47.01 |
| Noto-EQ-2024 | 36.49 | 37.33 | 36.56 | 38.35 | 42.84 | 40.57 | 45.78 |
| Average | 48.73 | 47.67 | 49.08 | 50.19 | 50.74 | 52.26 | 51.39 |
3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the proxy damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
4 Results and Analysis
4.1 Evaluation results on standard machine learning split
Table 5 shows the evaluation results for each model on the test set. We observe that ChangeMamba achieves the best overall performance, with an OA of 96.22, an mIoU of 67.63, and the highest and scores (90.90 and 72.70, respectively). DamageFormer also performs well, closely following ChangeMamba, with an mIoU of 67.09 and an OA of 96.13. Both models demonstrate strong capability in both building localization and damage classification tasks. The accuracy advantage of ChangeMamba and DamageFormer underscores the importance of leveraging advanced deep learning architectures to improve performance in complex tasks such as building damage assessment. For models using the direct prediction approach (UNet, DeepLabV3+, SiamAttnUNet), UNet delivers the best results, with an mIoU of 64.94 and an OA of 95.47. However, its performance still lags behind that of the decoupled models, further emphasizing the advantage of task decoupling.
To ensure that the evaluation is not dominated by a few events with a large number of images, e.g., Turkey-EQ-2023, Table 6 further presents the event-level mIoU for each model across 14 disaster events. ChangeMamba and DamageFormer achieve the highest average mIoU, with scores of 51.39 and 52.26, respectively. Notably, DamageFormer performs particularly well on events such as Beirut-EP-2020, Marshall-WF-2021, and Derna-FL-2023, demonstrating its robustness across different types of disasters. While performance varies across events, earthquake-related events such as Les Cayes-EQ-2021, Morocco-EQ-2023, and Noto-EQ-2024 present greater challenges for all models, where the average mIoU remains relatively low. This highlights the need for further research to improve model robustness for earthquake damage assessment, where damage patterns are often more complex and diverse.
Finally, Figure 6 shows some of the building damage proxy maps obtained by the seven models on the test set.
4.2 What have models learned and what can they learn?
To better understand model behavior beyond performance metrics, we explore the internal attention patterns of trained models using class activation maps (CAMs). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events, including volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. In Goma-VE-2021, for the “Destroyed” category, the encoder processing the optical imagery exhibits strong activations across nearly all built-up regions, accurately localizing individual buildings. This suggests the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, the encoder handling the SAR imagery shows intense activation over the lava-covered region on the left, indicating that the model has identified this as a key signal for destruction—likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, however, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder stage, the destroyed buildings in the lava-affected area are strongly activated, aligning well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as the subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
Beyond understanding what the models have already learned, a more critical question is: What can they learn? More specifically, to what extent can optical–SAR combinations support accurate building damage assessment across diverse disaster scenarios? To investigate this, we aggregate the mIoU of the seven models and show their average mIoU across seven major disaster types in Figure 8. Most notably, wildfire and volcano events exhibit the highest IoU scores for the “Destroyed” category, both exceeding 70. This indicates that the model can effectively leverage SAR-based backscatter anomalies, such as lava flows or widespread debris fields, to detect fully destroyed structures. These results demonstrate the strong potential of Optical–SAR fusion in such high-impact scenarios. However, for the “Damaged” category, the performance drops significantly. The average IoU for damaged buildings in wildfire events remains below 20, while in volcano events, it falls to around 5. This suggests that single-polarization SAR imagery lacks the fine-grained information needed to reliably distinguish partially damaged buildings, where structural integrity may still be partially preserved and backscatter signals remain ambiguous. For earthquake events, both damaged and destroyed categories yield relatively low IoUs. This is likely due to the complex and heterogeneous patterns of structural collapse typical of seismic events, where damage is often subtle, partial, and highly variable. These conditions pose significant challenges for SAR-based assessment. Interestingly, for flood and hurricane events, the model achieves relatively high IoU scores for the “Damaged” category with about 50 and 60, highlighting that SAR is effective at capturing contextual environmental changes—such as water inundation or terrain disruption—which indirectly aid in assessing building damage. In the case of the conflict event, the model’s performance on the destroyed class is surprisingly low. We attribute this to the limited number of destroyed samples in the dataset for this category, which leads to insufficient learning and poor generalization.
In summary, these findings confirm both the promise and limitations of Optical–SAR combinations for all-weather, global-scale disaster response. While the combination performs well in events characterized by large-scale surface disruption (e.g., wildfires, volcanoes), it struggles with subtle or localized damage patterns. Incorporating richer data sources, such as fully polarimetric SAR and LiDAR, can further enhance the accuracy and reliability of future building damage assessments.
| Method | mIoU (set) | mIoU (event) |
| Baseline | 67.63 | 51.39 |
| Test-time augmentation | 68.50 0.87 | 51.95 0.56 |
| Object-based major voting | 67.22 0.41 | 52.08 0.69 |
| Ensembling multiple models | 68.45 0.82 | 52.14 0.75 |
| All | 68.86 1.23 | 52.31 0.92 |
4.3 Effect of post-processing methods
Post-processing techniques help refine the raw predictions from deep learning models, reducing noise, improving consistency, and ensuring spatial coherence in damage proxy maps (Zheng et al., 2021). Here, we further explore the impact of post-processing algorithms. Table 7 presents the effect of three post-processing techniques applied to ChangeMamba. The post-processing methods evaluated include test-time augmentation, object-based major voting, and model ensembling. Further details on these methods are provided in Appendix D
Firstly, test-time augmentation improves the mIoU by 0.87 at the set level and 0.56 at the event level, demonstrating its effectiveness in enhancing model robustness across diverse disaster scenarios. Object-based major voting, which aggregates predictions at the building-object level to enforce spatial consistency, slightly reduces set-level mIoU (-0.41) but improves event-level mIoU (+0.69). Ensembling multiple models leads to a 0.82 increase in mIoU at the set level and a 0.75 increase at the event level, reinforcing its effectiveness in improving model performance across different disaster events. Applying all post-processing techniques together yields the highest performance improvement, with a 1.23 increase in set-level mIoU and a 0.92 increase in event-level mIoU. These results confirm that a combination of different post-processing methods can significantly enhance the reliability of AI-driven damage assessments, ensuring better generalization across disaster types and locations.
In summary, post-processing techniques play a crucial role in improving damage proxy maps. Future work can further explore adaptive post-processing strategies tailored to specific disaster types to enhance prediction reliability in multimodal EO data contexts.
4.4 Evaluation results on cross-event transfer split
4.4.1 Baseline methods
| Events | UNet | DeepLabV3+ | SiamAttnUNet | SiamCRNN | ChangeOS | DamageFormer | ChangeMamba |
| zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | |
| Beirut-EP-2020 | 31.62 / 34.89 | 34.72 / 33.58 | 35.33 / 29.17 | 39.43 / 40.94 | 38.48 / 38.21 | 38.71 / 39.93 | 40.30 / 40.92 |
| Bata-EP-2021 | 33.48 / 37.48 | 33.55 / 30.23 | 33.78 / 35.24 | 37.46 / 40.06 | 38.00 / 38.18 | 38.91 / 39.42 | 39.17 / 40.31 |
| Goma-VE-2021 | 55.97 / 57.14 | 55.01 / 58.13 | 57.52 / 55.75 | 58.91 / 59.54 | 58.66 / 57.85 | 55.86 / 59.89 | 60.39 / 61.91 |
| Les Cayes-EQ-2021 | 26.21 / 38.72 | 35.50 / 37.90 | 37.22 / 38.76 | 39.45 / 40.28 | 39.36 / 40.26 | 40.41 / 41.07 | 41.35 / 42.06 |
| La Palma-VE-2021 | 32.74 / 35.29 | 30.86 / 36.25 | 34.00 / 34.99 | 32.90 / 33.32 | 29.96 / 32.60 | 32.28 / 37.62 | 34.18 / 34.93 |
| Marshall-WF-2021 | 34.97 / 37.84 | 35.05 / 37.12 | 35.23 / 39.54 | 39.82 / 46.83 | 36.31 / 43.72 | 40.93 / 51.15 | 40.61 / 45.96 |
| Ukraine-AC-2022 | 32.36 / 37.50 | 32.23 / 35.98 | 33.84 / 35.76 | 34.77 / 38.14 | 36.30 / 39.28 | 33.84 / 36.51 | 35.78 / 36.99 |
| Turkey-EQ-2023 | 37.82 / 35.83 | 37.28 / 40.49 | 36.09 / 37.49 | 38.32 / 40.11 | 38.45 / 39.51 | 37.20 / 40.27 | 41.28 / 42.80 |
| Kyaukpyu-CC-2023 | 31.46 / 31.71 | 34.01 / 34.75 | 33.77 / 35.93 | 34.19 / 37.71 | 36.66 / 36.25 | 32.99 / 33.54 | 33.85 / 36.08 |
| Hawaii-WF-2023 | 31.73 / 39.80 | 31.32 / 45.44 | 31.79 / 45.12 | 30.94 / 42.65 | 31.15 / 41.96 | 31.60 / 50.64 | 36.14 / 50.89 |
| Morocco-EQ-2023 | 38.24 / 39.81 | 40.48 / 41.23 | 40.78 / 40.69 | 40.89 / 42.89 | 41.40 / 41.89 | 41.89 / 43.42 | 42.86 / 43.51 |
| Derna-FL-2023 | 33.25 / 38.44 | 35.07 / 40.02 | 33.76 / 42.15 | 35.88 / 44.77 | 35.80 / 43.85 | 36.27 / 44.69 | 37.02 / 45.47 |
| Acapulco-HC-2023 | 26.99 / 38.76 | 27.78 / 33.24 | 30.93 / 38.94 | 26.70 / 34.53 | 27.06 / 34.50 | 27.48 / 32.87 | 28.45 / 38.96 |
| Noto-EQ-2024 | 31.02 / 33.83 | 35.57 / 39.17 | 35.87 / 36.22 | 36.04 / 38.52 | 39.14 / 39.15 | 36.37 / 43.93 | 39.90 / 44.33 |
| Average | 34.13 / 38.36 | 35.60 / 38.82 | 36.42 / 38.98 | 37.55 / 41.45 | 37.62 / 40.52 | 37.48 / 42.49 | 39.38 / 43.23 |
Cross-event transfer, especially under zero-shot settings, poses a significant challenge for building damage assessment. As shown in Table 8, the average mIoU across all baseline models in the zero-shot setting is below 40, a noticeable drop compared to their performance under the fully supervised standard ML split, where models typically achieve 48 to 52 mIoU. This performance gap underscores the difficulty of generalizing to unseen disaster events without access to any target-domain supervision due to substantial domain shifts in imaging conditions, damage patterns, urban morphology, and sensor response. Despite this, all models exhibit clear performance gains in the one-shot setting, where a small number of labeled samples from the target event are available. This suggests that even minimal supervision can significantly aid adaptation to new disaster contexts.
Among the evaluated models, ChangeMamba consistently achieves the highest overall performance, with an average mIoU of 39.38 in zero-shot and 43.23 in one-shot settings. It is closely followed by DamageFormer, highlighting the strength of recent transformer- and Mamba-based architectures in transferring learned knowledge under complex multimodal and disaster scenarios. Models using a decoupled architecture generally outperform the direct segmentation models, confirming that separating building localization and damage classification improves generalization in cross-event transfer tasks. Conversely, events like Acapulco-HC-2023 and Ukraine-AC-2022 show weaker results across all models, reflecting the difficulty in transferring to domains with limited or inconsistent destruction patterns.
4.4.2 Unsupervised domain adaptation and semi-supervised learning methods
| Method | Avg. mIoU |
| Source-Only | 35.60 |
| AdaptSeg (Tsai et al., 2018) | 36.05 0.45 |
| AdvEnt (Vu et al., 2019) | 35.43 0.17 |
| CLAN (Luo et al., 2019) | 36.01 0.41 |
| PyCDA (Lian et al., 2019) | 33.19 2.41 |
| FDA (Yang and Soatto, 2020) | 36.29 0.69 |
| Method | Avg. mIoU |
| Baseline | 38.82 |
| MT (Tarvainen and Valpola, 2017) | 40.00 1.18 |
| CCT (Ouali et al., 2020) | 39.49 0.67 |
| GCT (Ke et al., 2020) | 39.34 0.52 |
| CPS (Chen et al., 2021) | 37.06 1.76 |
In the context of cross-event transfer, UDA and SSL methods naturally emerge as promising strategies to bridge the domain gap between source and target disaster events. In the zero-shot setting, UDA methods aim to improve model generalization by aligning the source and target domains without requiring any target labels. In the one-shot setting, SSL methods further leverage a small number of labeled samples along with abundant unlabeled data from the target event, making them especially appealing for real-world disaster scenarios where rapid annotation is infeasible. However, it remains an open question whether these UDA and SSL methods, originally developed for natural image domains, can effectively handle the challenges of multimodal EO data in complex disaster scenarios. To examine the applicability of these methods, we evaluate several representative UDA and SSL algorithms using DeepLabV3+ as the baseline model. The results are reported in Tables 9 and 10, respectively.
Table 9 reports the performance of five classical UDA methods on the BRIGHT dataset. Compared to the source-only baseline (35.60 mIoU), most methods achieve modest improvements, with FDA (+0.69) and AdaptSeg (+0.45) showing the most consistent gains. CLAN also performs slightly better (+0.41), suggesting that category-level alignment contributes positively even under a large domain shift. In contrast, AdvEnt shows negligible change, and PyCDA significantly underperforms, dropping over 2 mIoU compared to the baseline. This performance degradation indicates that approaches relying on pseudo-label refinement or curriculum learning may struggle in disaster imagery, where spatial layout and damage semantics vary drastically across events. Overall, while UDA methods show some promise, their improvements are relatively minor and not robust across all events, underscoring the difficulty of domain alignment in EO-based damage mapping.
In the one-shot setting, we evaluate several popular SSL methods, summarized in Table 10. With only minimal supervision in the target event, all methods except CPS achieve consistent improvements over the supervised one-shot baseline (38.82). The best-performing method is Mean Teacher (MT), which yields a gain of +1.18, followed by CCT (+0.67) and GCT (+0.52). These results show that simple consistency-based teacher-student frameworks are particularly effective in leveraging unlabeled data under limited supervision, likely due to their robustness to noisy or class-imbalanced targets.
4.4.3 Why cross-event transfer is challenging?
To better understand why cross-event generalization is so difficult, we explore two fundamental factors rooted in the nature of real-world disaster response.
1.
Inconsistent damage signatures across events. Figure 9 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. Firstly, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are not consistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. In addition, the distribution shift becomes more pronounced across different disaster types (the second row). For instance, the destroyed-building signatures in wildfires are notably different from those in volcano eruptions, which again differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster may learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
Figure 10: Comparison of models’ best performance (mIoU) on test events versus the best checkpoints selected on validation sets under the cross-event transfer setting. The farther a point lies from the diagonal line, the larger the gap between the model’s selected performance and its true upper bound. -
2.
Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target-domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a target domain validation set is available to tune hyperparameters or select the best checkpoint. However, in disaster response scenarios, no such validation data can be assumed. Figure 10 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source-domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source-domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future solutions should consider both aspects—developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).
4.5 Extended evaluation of unsupervised multimodal change detection methods
| Method | UMCD benchmark datasets | BRIGHT | |||||
| OA | F1 | KC | OA | F1 | IoU | KC | |
| IRG-McS (Sun et al., 2021) | 98.3 / - / 97.1 / 97.2 | 80.4 / - / 75.4 / 73.7 | 79.4 / - / 73.9 / 75.1 | 90.03 | 12.65 | 6.75 | 7.65 |
| SR-GCAE (Chen et al., 2022b) | 98.6 / 98.5 / - / - | 82.9 / 77.6 / - / - | 82.1 / 76.9 / - / - | 77.83 | 14.35 | 7.73 | 5.64 |
| FD-MCD (Chen et al., 2023) | 98.2 / 97.8 / - / 96.7 | 81.4 / 72.2 / - / 73.2 | 82.3 / 71.1 / - / 71.4 | 80.96 | 15.84 | 8.60 | 7.94 |
| AOSG (Han et al., 2024) | - / - / - / 96.4 | - / - / - / 77.7 | - / - / - / 75.9 | 77.93 | 10.75 | 5.68 | 3.98 |
| AGSCC (Sun et al., 2024a) | 98.3 / - / 95.9 / 97.6 | 78.2 / - / 68.0 / 77.9 | 77.3 / - / 65.8 / 76.6 | 88.49 | 14.82 | 8.00 | 9.54 |
| AEKAN (Liu et al., 2025) | 98.7 / - / - / 98.3 | 83.8 / - / - / 84.7 | 83.1/ - / - / 83.9 | 81.60 | 13.09 | 7.00 | 3.56 |
Unsupervised multimodal change detection (UMCD) plays a crucial role in post-disaster assessment by enabling rapid analysis of affected areas without requiring labeled data or prior model training (Lv et al., 2022). The current research is mainly conducted on toy datasets with limited geographic diversity and scene complexity. These datasets mainly involve general land-cover changes. It is still unknown whether these methods can work in real disaster occurrence/building damage scenarios. BRIGHT offers a new opportunity to evaluate UMCD methods, thereby providing insights into their robustness and scalability. Six representative UMCD methods (Sun et al., 2021; Chen et al., 2022b, 2023; Han et al., 2024; Sun et al., 2024a; Liu et al., 2025) are evaluated in this paper. The specific experimental setup is described in Appendix G.
Table 11 reports the performance of these methods on BRIGHT and their accuracy on current UMCD benchmark datasets shown in Figure 12. Although these methods can achieve considerable performance on the current benchmark dataset, their performance suffers noticeable declines on BRIGHT. For instance, while these methods achieve F1 scores between 70-85 on existing benchmarks, their F1 scores drop below 20 on BRIGHT. This dramatic performance gap underscores the limitations of current UMCD research and highlights the challenges posed by real disaster scenarios. We identify three primary reasons for this decline:
1.
Limitations in traditional UMCD datasets: The traditional UMCD datasets consist of only a handful of image pairs, often depicting simple land-cover changes such as urban expansion or deforestation, with relatively low-resolution imagery with limited geographic diversity, as shown in Figure 12. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, BRIGHT provides thousands of high-resolution multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
2.
Interference from non-disaster changes: Unlike previous UMCD benchmarks where general land-cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, BRIGHT requires methods to differentiate true building structural damage from irrelevant land-cover variations, posing a unique challenge to current approaches.
3.
Problematic evaluation protocol: Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real disaster scenario, there are no labeled samples available for hyperparameter tuning, making the current practice unrealistic. BRIGHT exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results highlight the challenges of applying existing UMCD models to real-world disaster scenarios. The BRIGHT dataset reveals significant limitations in current methodologies and also presents new opportunities for future UMCD research.
4.6 Extended evaluation of unsupervised multimodal image matching methods
| Method | Before (meters) | After (meters) |
| FLSS (Ye et al., 2017) | 16.97 | 10.01 |
| HOPC (Ye et al., 2019) | 9.37 | |
| LNIFT (Li et al., 2022) | N/A | |
| SRIF (Li et al., 2023a) | N/A |
Finally, precise image alignment is a critical prerequisite for any multimodal EO application. BRIGHT offers a unique opportunity to evaluate the performance of existing UMIM algorithms under realistic, large-scale disaster conditions. Due to the lack of real pixel-level ground truth correspondences, we adopt a proxy evaluation strategy using manually selected control points (as shown in Figure 11) as references. These points are selected by several EO experts to represent clearly identifiable and stable features across modalities. While this does not constitute absolute ground truth, the use of such human-verified correspondences provides a valuable reference, allowing us to assess how closely automated methods approximate human matching ability under multimodal and disaster conditions.
Table 12 presents the quantitative results of several representative UMIM methods on the Noto-Earthquake-2024 scene. Our experimental results reveal that traditional feature-based methods generally fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable. In contrast, area-based methods, which rely on region-level similarity rather than point-level correspondence, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that the BRIGHT benchmark and its challenging real-world settings will inspire the development of new UMIM methods that are more robust to large scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
5 Discussion
5.1 Limitation of BRIGHT
The performance of ML/DL models heavily depends on the quality of the training data. The BRIGHT dataset is designed to enable ML/DL models trained on it to generate building damage proxy maps of practical significance for future disaster events. It is anticipated that BRIGHT will serve as a benchmark for numerous subsequent studies and practical disaster relief applications.
However, the BRIGHT dataset also has some limitations. Firstly, our dataset involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 4. Secondly, the building polygons in BRIGHT were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on BRIGHT. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise. Finally, although BRIGHT is globally distributed and richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 v.s. Hawaii-Wildfire-2023 in Table 2. Although we used an additional event-wise evaluation method, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. It may affect the generalizability of the trained model in real disaster scenarios.
Overall, despite the limitations, it is the first time such a multimodal VHR dataset has been constructed that caters to a transferable approach to obtain building damage proxy maps at different geographical locations.
5.2 Significance of BRIGHT
Latency in responding to disaster events often arises from delays in interpretation methods and the acquisition of EO data (Ye et al., 2024). Traditionally, building damage proxy maps are generated through expert visual interpretation, a time-consuming process. Although ML- and DL-based methods offer automated and efficient approaches to building damage proxy mapping, they are often limited by the available training data. Large open-source optical datasets, though valuable, restrict models to optical imagery alone. The acquisition of clean optical EO data also requires sunlight and a cloud-free environment, which can result in delays of several days for certain disaster events such as typhoons and wildfires. As the first globally distributed multimodal dataset, BRIGHT encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on BRIGHT to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, BRIGHT offers several distinct characteristics: multimodal data, VHR imagery with resolutions better than 1 meter per pixel, coverage of six types of natural disasters, as well as human-made disasters, rich geographic diversity, and open access to the community. Due to these features, BRIGHT is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications. BRIGHT facilitates research on building damage assessment. Looking forward, BRIGHT can be expanded to include additional data modalities, such as height data, which would provide valuable topographic information to enhance disaster response strategies. Incorporating more bands in optical and SAR data could also broaden the applicability of the dataset.
In addition to its primary application in building damage assessment, the BRIGHT dataset holds significant potential for other domains, particularly in multimodal EO data-related research. Firstly, BRIGHT can be directly used as a building detection benchmark dataset with a rich geographic diversity and building patterns. Beyond that, researchers can also leverage BRIGHT to generate new datasets tailored to specific needs to explore new techniques including, but not limited to, image registration, land cover change detection, land cover mapping, height estimation, EO-driven visual question answering (VQA), improving the performance and reliability of models in complex environments. Moreover, BRIGHT is a potential resource for the development of remote sensing foundation models (Wang et al., 2023; Hong et al., 2024). These emerging models, designed to generalize across various tasks and datasets, benefit immensely from the rich multimodal data and extensive geographic diversity offered by BRIGHT. By incorporating both optical and SAR data, BRIGHT enables the pre-training of more robust foundation models capable of handling the complexities of real-world remote sensing applications. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024). We envision that BRIGHT, true to its name, will bring even a glimmer of bright to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
A prompt and comprehensive understanding of building damage following a disaster is critical for effective disaster response and relief operations. Currently, large-scale optical datasets support the development of automated ML/DL methods, but their use is limited by weather conditions and daylight availability. In contrast, all-weather disaster response can be facilitated by leveraging multimodal EO data. In this paper, we introduced BRIGHT, the first globally distributed, multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. BRIGHT includes pre-event optical imagery and post-event SAR imagery with very high spatial resolutions (better than 1 meter). Alongside the dataset, we evaluated several state-of-the-art models on BRIGHT. The experimental results demonstrate the effectiveness of BRIGHT in mapping building damage across unseen events. These findings provide both performance baselines and valuable insights into model design for future research. BRIGHT is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve BRIGHT’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.
The BRIGHT dataset and the code for training and testing benchmark methods (including code related to IEEE GRSS DFC 2025) are available at https://github.com/ChenHongruixuan/BRIGHT.
Appendix A Feature visualization
For the feature visualisation of buildings in different events in Figure 5c, we specifically employ DINOv2 (Oquab et al., 2024) to extract the high-dimensional features of buildings in different events. The corresponding reference maps are used to exclude background features. The features are then visualized using the t-SNE algorithm (van der Maaten and Hinton, 2008).
To visualize the spatial focus of models during building damage mapping, we apply Grad-CAM (Selvaraju et al., 2017), a gradient-based technique for generating class-specific activation maps. We perform Grad-CAM analysis on the ChangeOS model. Specifically, we compute the activation responses for the “Damaged” and “Destroyed” categories from the following components: the third or fourth ResNet block of the two branches of encoder, capturing mid- to high-level modality-specific features; the fusion module of the damage classification branch in the decoder, reflecting early-stage fusion and decision-making. These activations help reveal which spatial regions the model relies on when assessing different damage levels and how it utilizes information from optical and SAR modalities differently depending on the disaster context.
Appendix B Manual registration and estimating registration errors
Despite multiple rounds of meticulous registration and cross-validation of BRIGHT by several EO experts, registration errors between optical and SAR images could not be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we adopt a proxy method for registration error estimation:
1.
Feature Point Selection: Select many feature points (e.g., 3000 points) from both the optical and SAR images using a keypoint detection algorithm.
- 2.
3.
Feature Matching: Perform feature matching between the extracted feature points across the optical and SAR images.
4.
Registration Error Computation: After completing the matching process, compute the pixel distances between the registered point pairs. These distances serve as a proxy for registration error.
5.
Outlier Removal: To improve the reliability of the estimation, we exclude matched points with a root mean square error (RMSE) greater than a certain threshold (e.g., 20 pixels). This threshold is set because after multiple rounds of expert corrections and cross-checking, large registration errors have already been eliminated. Points exceeding this threshold can be considered outliers.
| Event | RMSE (All) | RMSE (Building) |
| Beirut-EP-2020 | 1.365 | 1.365 |
| Bata-EP-2021 | 1.226 | 1.160 |
| Goma-VE-2021 | 0.966 | 0.975 |
| Les Cayes-EQ-2021 | 1.392 | 1.377 |
| La Palma-VE-2021 | 0.976 | 0.944 |
| Marshall-WF-2021 | 1.263 | 1.266 |
| Ukraine-AC-2022 | 0.899 | 0.882 |
| Turkey-EQ-2023 | 0.937 | 0.901 |
| Kyaukpyu-CC-2023 | 0.970 | 0.998 |
| Hawaii-WF-2023 | 1.124 | 1.124 |
| Morocco-EQ-2023 | 1.088 | 1.050 |
| Derna-FL-2023 | 1.267 | 1.255 |
| Acapulco-HC-2023 | 1.166 | 1.148 |
| Noto-EQ-2024 | 1.072 | 1.043 |
By following this method, we can estimate the registration errors between optical and SAR images in the dataset, providing a reliable foundation for further damage assessment analysis.
Figure of selected feature points (can be used as proxy ground truth for evaluating descriptors)
Appendix C Details of benchmark DL models
Here, we present some implementation details of our baseline model.
For UNet, the encoder has five convolutional blocks, each consisting of two 33 convolutional layers and two batch normalization layers. The number of channels in the five blocks are 64, 128, 256, 512, and 1024, respectively. The decoder contains four convolutional blocks, and the number of channels is set accordingly to the number of channels of the skip-connected features extracted by the encoder.
For DeepLabV3+, we applied ResNet-50 initialized with ImageNet pretraining weights (He et al., 2016) as the encoder. We then modified the input channel of the encoder so that the network can predict damage proxy maps directly from the optical and SAR stacked images.
For SiamAttnUNet, we followed the settings in its original literature (Adriano et al., 2021).
For SiamCRNN, we applied pseudo-siamese ResNet-18 as encoder and four ConvLSTM layers (Shi et al., 2015) with 33 convolutional kernel size and a hidden dimension of 128, as the decoder.
For ChangeOS, we used the ResNet-18 initialized with ImageNet pretraining weights as the encoder. We modified the encoder to a pseudo-siamese structure to extract features from pre-event and post-event images with different modalities.
For the ChangeMamba family of models (Chen et al., 2024), we chose one of them, MambaBDA-Tiny, for our experiments. We modified the encoder to a pseudo-siamese structure to extract features from pre-event and post-event images with different modalities.
Appendix D Details of post-processing methods
To further refine the raw predictions from deep learning models and enhance the quality of damage proxy maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 4.3. This section provides an overview of these methods and details the specific implementation used in our study.
Test-time augmentation is a widely used technique to improve model robustness by applying transformations to the input images at inference time. Instead of making a single prediction per image, multiple augmented versions of the same input are passed through the network, and the resulting predictions are aggregated. This reduces model sensitivity to spatial variations and increases prediction stability. In our study, we apply rotation (90°, 180°, 270°) and horizontal/vertical flipping to the input images during inference. The network produces a set of predictions for each augmented version, which are then aggregated at the logit level by summing the outputs before applying the softmax function.
Object-based post-processing aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
Model ensembling is a technique that combines predictions from multiple models to reduce uncertainty and improve generalization. By leveraging diverse models with different architectures, ensembling helps smooth out individual model biases and enhances the overall robustness of predictions. We perform ensembling by combining the results of three top-performing models: ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). Each model independently generates a damage proxy map, and the final prediction is obtained by averaging their logits before applying the softmax function.
Appendix E Details of unsupervised domain adaptation methods
| Events | AdaptSeg | AdvEnt | CLAN | PyCDA | FDA |
| Beirut-EP-2020 | 35.55 | 33.07 | 37.25 | 32.15 | 35.89 |
| Bata-EP-2021 | 37.54 | 35.20 | 36.98 | 30.01 | 36.31 |
| Goma-VE-2021 | 52.49 | 51.54 | 56.91 | 49.67 | 52.46 |
| Les Cayes-EQ-2021 | 37.11 | 36.86 | 39.47 | 33.40 | 33.90 |
| La Palma-VE-2021 | 30.54 | 30.04 | 30.10 | 32.76 | 30.50 |
| Marshall-WF-2021 | 37.20 | 35.01 | 34.35 | 30.14 | 33.93 |
| Ukraine-AC-2022 | 32.36 | 32.65 | 32.21 | 32.92 | 35.03 |
| Turkey-EQ-2023 | 33.69 | 34.54 | 36.68 | 34.58 | 36.82 |
| Kyaukpyu-CC-2023 | 34.62 | 29.22 | 30.83 | 31.31 | 36.71 |
| Hawaii-WF-2023 | 36.52 | 42.17 | 33.36 | 34.30 | 36.19 |
| Morocco-EQ-2023 | 37.93 | 37.42 | 39.04 | 30.55 | 38.76 |
| Derna-FL-2023 | 36.69 | 35.31 | 35.24 | 37.31 | 38.53 |
| Acapulco-HC-2023 | 27.05 | 27.32 | 27.73 | 30.11 | 31.68 |
| Noto-EQ-2024 | 35.49 | 35.65 | 33.99 | 25.44 | 31.39 |
| Average | 36.05 | 35.43 | 36.01 | 33.19 | 36.29 |
To evaluate the effectiveness of unsupervised domain adaptation (UDA) techniques in the cross-event zero-shot setting, we selected several representative methods from the computer vision literature originally designed for semantic segmentation tasks. The following methods are implemented:
AdaptSeg (Tsai et al., 2018): an adversarial training framework that aligns output space distributions between source and target domains.
AdvEnt (Vu et al., 2019): a variant of adversarial adaptation focusing on entropy minimization and structured output alignment.
CLAN (Luo et al., 2019): a category-level alignment method that selectively aligns class-specific features.
PyCDA (Lian et al., 2019): a curriculum domain adaptation framework that iteratively refines pseudo-labels across domains.
FDA (Yang and Soatto, 2020): a simple style-transfer-based approach that transfers source images into the target domain appearance via frequency mixing.
All methods are implemented using DeepLabV3+ as the backbone architecture. This choice is motivated by the fact that DeepLab-based networks are commonly used in the official implementations of these UDA methods, and many codebases are tightly coupled with DeepLab’s structure. Therefore, adopting DeepLabV3+ as the shared backbone ensures compatibility with existing implementations and avoids the need for extensive code modification. It also provides a fair and consistent basis for comparison across all methods.
In our experimental setup, each test disaster event is treated as the target domain, and only its unlabeled optical-SAR image pairs are used for adaptation. Importantly, no labeled target domain data is used for training or model selection, in line with our zero-shot assumption. This setup is fundamentally different from conventional UDA protocols in the computer vision community, where a target domain validation set is typically available to select the best model checkpoint, and the target test set is held out purely for evaluation. In contrast, our setting is designed to reflect real-world disaster response scenarios, where it is realistic to obtain unlabeled imagery, but no ground-truth annotations are available at adaptation time. Therefore, model selection must rely solely on source-domain feedback, which introduces additional challenges and more closely aligns with the operational constraints of emergency response.
The hyperparameters for each method closely follow the original papers to preserve method fidelity. Below are additional implementation details specific to certain methods:
FDA: For each training iteration, we generate frequency-transferred source images using FDA and mix them 1:1 with the original source images. The combined dataset is used to train the model from scratch.
PyCDA: Training proceeds in two stages. First, the model is trained on the source domain for 30,000 iterations. In the second stage, the adaptation module is applied to the target domain using pseudo-labeling and uncertainty-aware refinement as proposed in the original work.
For all methods, the same data preprocessing pipeline and training schedule (e.g., learning rate, batch size) are used unless otherwise specified in the original method, like those supervised models. The final evaluation is conducted on the target domain’s test set, and performance is reported in terms of mIoU.
Appendix F Details of semi-supervised learning methods
| Events | MT | CCT | GCT | CPS |
| Beirut-EP-2020 | 32.91 | 35.68 | 35.48 | 35.46 |
| Bata-EP-2021 | 37.94 | 36.61 | 38.25 | 36.96 |
| Goma-VE-2021 | 56.46 | 57.68 | 53.52 | 47.49 |
| Les Cayes-EQ-2021 | 40.20 | 39.57 | 38.83 | 39.41 |
| La Palma-VE-2021 | 33.52 | 33.03 | 36.90 | 33.29 |
| Marshall-WF-2021 | 39.73 | 40.12 | 38.74 | 32.94 |
| Ukraine-AC-2022 | 39.11 | 36.34 | 36.44 | 31.34 |
| Turkey-EQ-2023 | 38.96 | 39.72 | 36.50 | 37.15 |
| Kyaukpyu-CC-2023 | 36.72 | 35.84 | 33.56 | 30.22 |
| Hawaii-WF-2023 | 54.10 | 48.93 | 49.99 | 47.64 |
| Morocco-EQ-2023 | 37.18 | 40.54 | 41.10 | 37.54 |
| Derna-FL-2023 | 37.55 | 41.07 | 40.45 | 40.28 |
| Acapulco-HC-2023 | 32.30 | 30.16 | 33.36 | 31.51 |
| Noto-EQ-2024 | 43.40 | 37.61 | 37.74 | 37.65 |
| Average | 40.00 | 39.49 | 39.34 | 37.06 |
In the one-shot cross-event transfer setting, we evaluate several representative SSL methods for semantic segmentation. These methods were originally developed for natural image domains and are adapted here to the multimodal EO-based building damage assessment scenario.
Mean Teacher (MT) (Tarvainen and Valpola, 2017) maintains an exponential moving average (EMA) of model weights to generate stable predictions for unlabeled data, enforcing consistency between the student and teacher outputs.
CCT (Ouali et al., 2020) applies perturbations to the input data and enforces consistency between different views of the same image using multiple decoder branches.
GCT (Ke et al., 2020) further enhances this idea by introducing guidance from a pre-trained model to regularize learning.
CPS (Chen et al., 2021) trains two separate networks that generate pseudo-labels for each other, encouraging cross-supervised learning on unlabeled data.
All methods also use DeepLabV3+ as the backbone network. This decision is consistent with their original implementations and allows for seamless integration with official codebases without the need for architectural modifications. It also ensures a fair and consistent comparison across methods.
In the setup, 13 labeled disaster events serve as the training and validation sets, while a single labeled sample from the test (target) event is provided to simulate a realistic one-shot adaptation scenario. In addition, the remaining unlabeled samples from the test event are made available and used to facilitate SSL.
Each model is trained in two stages. In the first stage, the model is trained for 10,000 iterations using only standard supervised losses, including cross-entropy and Lovász-Softmax on the labeled data from the source domain and the single labeled sample from the target domain. This stage serves to initialize the model with a stable representation before semi-supervised training begins. In the second stage, each SSL method applies its respective semi-supervised objective using the unlabeled data from the target domain. These include consistency regularization (e.g., Mean Teacher), cross-view perturbation training (e.g., CCT and GCT), or dual-network pseudo supervision (e.g., CPS). Hyperparameters for each method follow the configurations suggested in their original publications.
Final performance is evaluated on the test set of the target domain using mIoU as the primary evaluation metric.
Appendix G Introduction of UMCD methods and a more practical evaluation protocol
We further evaluate several recent UMCD methods on the BRIGHT dataset. These methods are designed to detect binary changes from paired multimodal imagery without relying on manual labels.
IRG-McS (Sun et al., 2021) constructs superpixel-based graphs for each modality and iteratively matches their structural relationships to detect changes across heterogeneous image pairs, followed by Markov co-segmentation for final change map generation.
SR-GCAE (Chen et al., 2022b) learns robust graph-based representations of local and nonlocal structural relationships in multimodal images using a graph convolutional autoencoder, enabling effective change detection without supervision.
FD-MCD (Chen et al., 2023) proposes a Fourier domain framework that analyzes both local and nonlocal structural relationships using graph spectral convolution and adaptive fusion, enabling robust unsupervised multimodal change detection.
AOSG (Han et al., 2024) constructs an adaptively optimized structured graph to capture patch-level structural features in multimodal images, iteratively refining change intensity measures by fusing self-change and cross-domain structural differences for accurate unsupervised change detection.
AGSCC (Sun et al., 2024a) translates multimodal images via a structure cycle-consistent image regression framework that enforces similarity in structural graphs across domains, using adaptive graphs and multiple regularization terms to robustly detect changes without supervision.
AEKAN (Liu et al., 2025) utilizes a superpixel-based Siamese AutoEncoder built on Kolmogorov–Arnold Networks (KAN) to extract latent commonality features between multimodal images, using reconstruction and hierarchical consistency losses to detect changes in an unsupervised manner.
The training and evaluation protocols commonly adopted in the UMCD literature are often limited to individual image pairs (shown in Figure 12), using the same scene for representation learning, hyperparameter tuning, and evaluation. While this is understandable due to the lack of large-scale public datasets in this field, such a setup fails to reflect real-world use cases and often leads to overfitting and overestimated performance. To address this, we introduce a standardized and practical evaluation protocol using the BRIGHT dataset. Specifically, we use the validation set from the standard ML split in Section 2.3 as the training set for UMCD methods, including any hyperparameter or threshold tuning. These models are then evaluated on the combined training and test sets, which are strictly held out during adaptation. This avoids overlap between tuning and evaluation scenes and better simulates the real-world deployment setting, where a model is expected to generalize to new, unseen disasters.
For evaluation, it is important to note that most of these methods were originally developed for land cover change detection, where changes correspond to transitions between semantic categories such as vegetation, water, or built-up areas. Applying them directly to disaster damage detection poses certain challenges. In particular, “Damaged” buildings do not always exhibit strong spatial or spectral signals, making them hard to distinguish in an unsupervised setting. However, “Destroyed” buildings often result in a complete change of land cover appearance (e.g., collapsed structures, debris, or ground exposure), which aligns better with the assumptions of existing UMCD methods. Therefore, in our evaluation, we treat buildings labeled as “Destroyed” in BRIGHT as the positive (changed) class, and all other regions (including intact, damaged, and background) as the unchanged class. This allows a fair adaptation of these methods to a disaster-relevant setting while respecting their original design.
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (lead). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (equal). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (equal). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan999Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).
The contact author has declared that none of the authors has any competing interests.
Acknowledgements.
This work was supported in part by the JSPS, KAKENHI under Grant Number 24KJ0652 and 22H03609, the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S, Microsoft Research Asia, Next Generation AI Research Center of The University of Tokyo, and Young Researchers Exchange Programme between Japan and Switzerland under the Japanese-Swiss Science and Technology Programme. The authors would also like to give special thanks to Sarah Preston of Capella Space, Capella Space’s Open Data Gallery, Maxar Open Data Program, and Umbra Space’s Open Data Program for providing the valuable data.References
- Adriano et al. (2019) Adriano, B., Xia, J., Baier, G., Yokoya, N., and Koshimura, S.: Multi-Source Data Fusion Based on Ensemble Learning for Rapid Building Damage Mapping during the 2018 Sulawesi Earthquake and Tsunami in Palu, Indonesia, Remote Sensing, 11, 2019.
- Adriano et al. (2021) Adriano, B., Yokoya, N., Xia, J., Miura, H., Liu, W., Matsuoka, M., and Koshimura, S.: Learning from multimodal and multitemporal earth observation data for building damage mapping, ISPRS J. Photogramm. Remote Sens., 175, 132–143, 2021.
- Arciniegas et al. (2007) Arciniegas, G. A., Bijker, W., Kerle, N., and Tolpekin, V. A.: Coherence- and Amplitude-Based Analysis of Seismogenic Damage in Bam, Iran, Using ENVISAT ASAR Data, IEEE Transactions on Geoscience and Remote Sensing, 45, 1571–1581, 2007.
- Artés et al. (2019) Artés, T., Oom, D., de Rigo, D., Durrant, T. H., Maianti, P., Libertà, G., and San-Miguel-Ayanz, J.: A global wildfire dataset for the analysis of fire regimes and fire behaviour, Scientific Data, 6, 296, 2019.
- Bai et al. (2017) Bai, Y., Adriano, B., Mas, E., and Koshimura, S.: Machine Learning Based Building Damage Mapping from the ALOS-2/PALSAR-2 SAR Imagery: Case Study of 2016 Kumamoto Earthquake, Journal of Disaster Research, 12, 646–655, 2017.
- Bai et al. (2018) Bai, Y., Gao, C., Singh, S., Koch, M., Adriano, B., Mas, E., and Koshimura, S.: A Framework of Rapid Regional Tsunami Damage Recognition From Post-event TerraSAR-X Imagery Using Deep Neural Networks, IEEE Geoscience and Remote Sensing Letters, 15, 43–47, 2018.
- Bastos Moroz and Thieken (2024) Bastos Moroz, C. and Thieken, A. H.: Urban growth and spatial segregation increase disaster risk: lessons learned from the 2023 disaster on the North Coast of São Paulo, Brazil, Natural Hazards and Earth System Sciences, 24, 3299–3314, 2024.
- Berman et al. (2018) Berman, M., Triki, A. R., and Blaschko, M. B.: The Lovász-Softmax Loss: A Tractable Surrogate for the Optimization of the Intersection-Over-Union Measure in Neural Networks, in: Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition (CVPR), 2018.
- Bonafilia et al. (2020) Bonafilia, D., Tellman, B., Anderson, T., and Issenberg, E.: Sen1Floods11: A Georeferenced Dataset to Train and Test Deep Learning Flood Algorithms for Sentinel-1, in: Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) Workshops, 2020.
- Brett and Guida (2013) Brett, P. T. B. and Guida, R.: Earthquake Damage Detection in Urban Areas Using Curvilinear Features, IEEE Transactions on Geoscience and Remote Sensing, 51, 4877–4884, 2013.
- Brunner et al. (2010) Brunner, D., Lemoine, G., and Bruzzone, L.: Earthquake damage assessment of buildings using VHR optical and SAR imagery, IEEE Trans. Geosci. Remote Sens., 48, 2403–2420, 2010.
- Chen et al. (2020) Chen, H., Wu, C., Du, B., Zhang, L., and Wang, L.: Change Detection in Multisource VHR Images via Deep Siamese Convolutional Multiple-Layers Recurrent Neural Network, IEEE Trans. Geosci. Remote Sens., 58, 2848–2864, 2020.
- Chen et al. (2022a) Chen, H., Nemni, E., Vallecorsa, S., Li, X., Wu, C., and Bromley, L.: Dual-Tasks Siamese Transformer Framework for Building Damage Assessment, in: International Geoscience and Remote Sensing Symposium (IGARSS), pp. 1600–1603, 2022a.
- Chen et al. (2022b) Chen, H., Yokoya, N., Wu, C., and Du, B.: Unsupervised Multimodal Change Detection Based on Structural Relationship Graph Representation Learning, IEEE Trans. Geosci. Remote Sens., pp. 1–18, 2022b.
- Chen et al. (2023) Chen, H., Yokoya, N., and Chini, M.: Fourier domain structural relationship analysis for unsupervised multimodal change detection, ISPRS J. Photogramm. Remote Sens., 198, 99–114, 2023.
- Chen et al. (2024) Chen, H., Song, J., Han, C., Xia, J., and Yokoya, N.: ChangeMamba: Remote Sensing Change Detection With Spatiotemporal State Space Model, IEEE Transactions on Geoscience and Remote Sensing, 62, 1–20, 2024.
- Chen et al. (2018) Chen, L.-C., Zhu, Y., Papandreou, G., Schroff, F., and Adam, H.: Encoder-Decoder with Atrous Separable Convolution for Semantic Image Segmentation, in: Computer Vision – ECCV 2018, pp. 833–851, 2018.
- Chen and Sato (2013) Chen, S.-W. and Sato, M.: Tsunami Damage Investigation of Built-Up Areas Using Multitemporal Spaceborne Full Polarimetric SAR Images, IEEE Transactions on Geoscience and Remote Sensing, 51, 1985–1997, 2013.
- Chen et al. (2021) Chen, X., Yuan, Y., Zeng, G., and Wang, J.: Semi-Supervised Semantic Segmentation with Cross Pseudo Supervision, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 2613–2622, 2021.
- Cheng et al. (2021) Cheng, C.-S., Behzadan, A. H., and Noshadravan, A.: Deep learning for post-hurricane aerial damage assessment of buildings, Computer-Aided Civil and Infrastructure Engineering, 36, 695–710, https://doi.org/10.1111/mice.12658, 2021.
- Chini et al. (2009) Chini, M., Pierdicca, N., and Emery, W. J.: Exploiting SAR and VHR Optical Images to Quantify Damage Caused by the 2003 Bam Earthquake, IEEE Transactions on Geoscience and Remote Sensing, 47, 145–152, 2009.
- Chini et al. (2015) Chini, M., Anniballe, R., Bignami, C., Pierdicca, N., Mori, S., and Stramondo, S.: Identification of building double-bounces feature in very high resoultion SAR data for earthquake damage mapping, in: 2015 IEEE International Geoscience and Remote Sensing Symposium (IGARSS), pp. 2723–2726, 2015.
- Fan et al. (2019) Fan, X., Nie, G., Deng, Y., An, J., Zhou, J., Xia, C., and Pang, X.: Estimating earthquake-damage areas using Landsat-8 OLI surface reflectance data, International Journal of Disaster Risk Reduction, 33, 275–283, 2019.
- Freire et al. (2014) Freire, S., Santos, T., Navarro, A., Soares, F., Silva, J., Afonso, N., Fonseca, A., and Tenedório, J.: Introducing mapping standards in the quality assessment of buildings extracted from very high resolution satellite imagery, ISPRS Journal of Photogrammetry and Remote Sensing, 90, 1–9, 2014.
- Fujita et al. (2017) Fujita, A., Sakurada, K., Imaizumi, T., Ito, R., Hikosaka, S., and Nakamura, R.: Damage detection from aerial images via convolutional neural networks, in: 2017 Fifteenth IAPR International Conference on Machine Vision Applications (MVA), pp. 5–8, 2017.
- Ge et al. (2019) Ge, P., Gokon, H., Meguro, K., and Koshimura, S.: Study on the Intensity and Coherence Information of High-Resolution ALOS-2 SAR Images for Rapid Massive Landslide Mapping at a Pixel Level, Remote Sensing, 11, 2019.
- Ge et al. (2020) Ge, P., Gokon, H., and Meguro, K.: A review on synthetic aperture radar-based building damage assessment in disasters, Remote Sensing of Environment, 240, 111 693, 2020.
- Gerard et al. (2024) Gerard, S., Borne-Pons, P., and Sullivan, J.: A simple, strong baseline for building damage detection on the xBD dataset, arXiv preprint arXiv:2401.17271, 2024.
- Ghorbanzadeh et al. (2022) Ghorbanzadeh, O., Xu, Y., Ghamisi, P., Kopp, M., and Kreil, D.: Landslide4Sense: Reference Benchmark Data and Deep Learning Models for Landslide Detection, IEEE Transactions on Geoscience and Remote Sensing, 60, 1–17, 2022.
- Guo et al. (2024) Guo, H., Su, X., Wu, C., Du, B., and Zhang, L.: SAAN: Similarity-Aware Attention Flow Network for Change Detection With VHR Remote Sensing Images, IEEE Transactions on Image Processing, 33, 2599–2613, 2024.
- Gupta and Shah (2021) Gupta, R. and Shah, M.: RescueNet: Joint Building Segmentation and Damage Assessment from Satellite Imagery, in: 2020 25th International Conference on Pattern Recognition (ICPR), pp. 4405–4411, 2021.
- Gupta et al. (2019) Gupta, R., Goodman, B., Patel, N., Hosfelt, R., Sajeev, S., Heim, E., Doshi, J., Lucas, K., Choset, H., and Gaston, M.: Creating xBD: A Dataset for Assessing Building Damage from Satellite Imagery, in: Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) Workshops, 2019.
- Han et al. (2024) Han, T., Tang, Y., Zou, B., and Feng, H.: Unsupervised multimodal change detection based on adaptive optimization of structured graph, International Journal of Applied Earth Observation and Geoinformation, 126, 103 630, 2024.
- He et al. (2016) He, K., Zhang, X., Ren, S., and Sun, J.: Deep residual learning for image recognition, Proceedings of the IEEE Computer Society Conference on Computer Vision and Pattern Recognition (CVPR), pp. 770–778, 2016.
- He et al. (2024) He, K., Shen, X., and Anagnostou, E. N.: A global forest burn severity dataset from Landsat imagery (2003–2016), Earth System Science Data, 16, 3061–3081, 2024.
- Hong et al. (2024) Hong, D., Zhang, B., Li, X., Li, Y., Li, C., Yao, J., Yokoya, N., Li, H., Ghamisi, P., Jia, X., Plaza, A., Gamba, P., Benediktsson, J. A., and Chanussot, J.: SpectralGPT: Spectral Remote Sensing Foundation Model, IEEE Transactions on Pattern Analysis and Machine Intelligence, 46, 5227–5244, 2024.
- Huot et al. (2022) Huot, F., Hu, R. L., Goyal, N., Sankar, T., Ihme, M., and Chen, Y.-F.: Next Day Wildfire Spread: A Machine Learning Dataset to Predict Wildfire Spreading From Remote-Sensing Data, IEEE Transactions on Geoscience and Remote Sensing, 60, 1–13, 2022.
- Karimzadeh and Mastuoka (2017) Karimzadeh, S. and Mastuoka, M.: Building Damage Assessment Using Multisensor Dual-Polarized Synthetic Aperture Radar Data for the 2016 M 6.2 Amatrice Earthquake, Italy, Remote Sensing, 9, 2017.
- Kaur et al. (2023) Kaur, N., Lee, C.-C., Mostafavi, A., and Mahdavi-Amiri, A.: Large-scale building damage assessment using a novel hierarchical transformer architecture on satellite images, Computer-Aided Civil and Infrastructure Engineering, 38, 2072–2091, 2023.
- Ke et al. (2020) Ke, Z., Qiu, D., Li, K., Yan, Q., and Lau, R. W. H.: Guided Collaborative Training for Pixel-Wise Semi-Supervised Learning, in: European Conference Computer Vision, p. 429–445, 2020.
- Kreibich et al. (2022) Kreibich, H., Loon, A. F. V., Schröter, K., Ward, P. J., Mazzoleni, M., Sairam, N., Abeshu, G. W., Agafonova, S., AghaKouchak, A., Aksoy, H., Alvarez-Garreton, C., Aznar, B., Balkhi, L., Barendrecht, M. H., Biancamaria, S., Bos-Burgering, L., Bradley, C., Budiyono, Y., Buytaert, W., Capewell, L., Carlson, H., Cavus, Y., Couasnon, A., Coxon, G., Daliakopoulos, I., de Ruiter, M. C., Delus, C., Erfurt, M., Esposito, G., François, D., Frappart, F., Freer, J., Frolova, N., Gain, A. K., Grillakis, M., Grima, J. O., Guzmán, D. A., Huning, L. S., Ionita, M., Kharlamov, M., Khoi, D. N., Kieboom, N., Kireeva, M., Koutroulis, A., Lavado-Casimiro, W., Li, H.-Y., Llasat, M. C., Macdonald, D., Mård, J., Mathew-Richards, H., McKenzie, A., Mejia, A., Mendiondo, E. M., Mens, M., Mobini, S., Mohor, G. S., Nagavciuc, V., Ngo-Duc, T., Huynh, T. T. N., Nhi, P. T. T., Petrucci, O., Nguyen, H. Q., Quintana-Seguí, P., Razavi, S., Ridolfi, E., Riegel, J., Sadik, M. S., Savelli, E., Sazonov, A., Sharma, S., Sörensen, J., Souza, F. A. A., Stahl, K., Steinhausen, M., Stoelzle, M., Szalińska, W., Tang, Q., Tian, F., Tokarczyk, T., Tovar, C., Tran, T. V. T., Huijgevoort, M. H. J. V., van Vliet, M. T. H., Vorogushyn, S., Wagener, T., Wang, Y., Wendt, D. E., Wickham, E., Yang, L., Zambrano-Bigiarini, M., Blöschl, G., and Baldassarre, G. D.: The challenge of unprecedented floods and droughts in risk management, Nature, 608, 80–86, 2022.
- Li et al. (2024) Li, C., Hong, D., Zhang, B., Liao, T., Yokoya, N., Ghamisi, P., Chen, M., Wang, L., Benediktsson, J. A., and Chanussot, J.: Interpretable foundation models as decryptors peering into the Earth system, The Innovation, 5, 100 682, 2024.
- Li et al. (2020) Li, J., Hu, Q., and Ai, M.: RIFT: Multi-Modal Image Matching Based on Radiation-Variation Insensitive Feature Transform, IEEE Transactions on Image Processing, 29, 3296–3310, 2020.
- Li et al. (2022) Li, J., Xu, W., Shi, P., Zhang, Y., and Hu, Q.: LNIFT: Locally Normalized Image for Rotation Invariant Multimodal Feature Matching, IEEE Transactions on Geoscience and Remote Sensing, 60, 1–14, 2022.
- Li et al. (2023a) Li, J., Hu, Q., and Zhang, Y.: Multimodal image matching: A scale-invariant algorithm and an open dataset, ISPRS Journal of Photogrammetry and Remote Sensing, 204, 77–88, 2023a.
- Li et al. (2023b) Li, T., Wang, C., Zhang, H., Wu, F., and Zheng, X.: DDFormer: A Dual-Domain Transformer for Building Damage Detection Using High-Resolution SAR Imagery, IEEE Geoscience and Remote Sensing Letters, 20, 1–5, 2023b.
- Lian et al. (2019) Lian, Q., Duan, L., Lv, F., and Gong, B.: Constructing Self-Motivated Pyramid Curriculums for Cross-Domain Semantic Segmentation: A Non-Adversarial Approach, in: IEEE/CVF International Conference on Computer Vision (ICCV), pp. 6757–6766, 2019.
- Liu et al. (2025) Liu, T., Xu, J., Lei, T., Wang, Y., Du, X., Zhang, W., Lv, Z., and Gong, M.: AEKAN: Exploring Superpixel-Based AutoEncoder Kolmogorov-Arnold Network for Unsupervised Multimodal Change Detection, IEEE Transactions on Geoscience and Remote Sensing, 63, 1–14, 2025.
- Liu and Yamazaki (2017) Liu, W. and Yamazaki, F.: Extraction of Collapsed Buildings in the 2016 Kumamoto Earthquake Using Multi-Temporal PALSAR-2 Data, Journal of Disaster Research, 12, 241–250, 10.20965/jdr.2017.p0241, 2017.
- Liu et al. (2013) Liu, W., Yamazaki, F., Gokon, H., and ichi Koshimura, S.: Extraction of Tsunami-Flooded Areas and Damaged Buildings in the 2011 Tohoku-Oki Earthquake from TerraSAR-X Intensity Images, Earthquake Spectra, 29, 183–200, 2013.
- Liu et al. (2021) Liu, Z., Lin, Y., Cao, Y., Hu, H., Wei, Y., Zhang, Z., Lin, S., and Guo, B.: Swin Transformer: Hierarchical Vision Transformer using Shifted Windows, in: IEEE/CVF International Conference on Computer Vision (ICCV), pp. 9992–10 002, 2021.
- Loshchilov and Hutter (2017) Loshchilov, I. and Hutter, F.: Decoupled weight decay regularization, arXiv preprint arXiv:1711.05101, 2017.
- Luo et al. (2019) Luo, Y., Zheng, L., Guan, T., Yu, J., and Yang, Y.: Taking a Closer Look at Domain Shift: Category-Level Adversaries for Semantics Consistent Domain Adaptation, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 2502–2511, 2019.
- Lv et al. (2022) Lv, Z., Huang, H., Li, X., Zhao, M., Benediktsson, J. A., Sun, W., and Falco, N.: Land Cover Change Detection With Heterogeneous Remote Sensing Images: Review, Progress, and Perspective, Proceedings of the IEEE, 110, 1976–1991, 2022.
- Matsuoka and Yamazaki (2005) Matsuoka, M. and Yamazaki, F.: Building Damage Mapping of the 2003 Bam, Iran, Earthquake Using Envisat/ASAR Intensity Imagery, Earthquake Spectra, 21, 285–294, 2005.
- Matsuoka and Yamazaki (2010) Matsuoka, M. and Yamazaki, F.: Comparative analysis for detecting areas with building damage from several destructive earthquakes using satellite synthetic aperture radar images, Journal of Applied Remote Sensing, 4, 041 867, 2010.
- Matsuoka et al. (2010) Matsuoka, M., Koshimura, S., and Nojima, N.: Estimation of building damage ratio due to earthquakes and tsunamis using satellite SAR imagery, in: 2010 IEEE International Geoscience and Remote Sensing Symposium, pp. 3347–3349, 2010.
- Meena et al. (2023) Meena, S. R., Nava, L., Bhuyan, K., Puliero, S., Soares, L. P., Dias, H. C., Floris, M., and Catani, F.: HR-GLDD: a globally distributed dataset using generalized deep learning (DL) for rapid landslide mapping on high-resolution (HR) satellite imagery, Earth System Science Data, 15, 3283–3298, 2023.
- Nguyen et al. (2017) Nguyen, D. T., Ofli, F., Imran, M., and Mitra, P.: Damage Assessment from Social Media Imagery Data During Disasters, in: Proceedings of the 2017 IEEE/ACM International Conference on Advances in Social Networks Analysis and Mining 2017, p. 569–576, 2017.
- Oquab et al. (2024) Oquab, M., Darcet, T., Moutakanni, T., Vo, H. V., Szafraniec, M., Khalidov, V., Fernandez, P., HAZIZA, D., Massa, F., El-Nouby, A., Assran, M., Ballas, N., Galuba, W., Howes, R., Huang, P.-Y., Li, S.-W., Misra, I., Rabbat, M., Sharma, V., Synnaeve, G., Xu, H., Jegou, H., Mairal, J., Labatut, P., Joulin, A., and Bojanowski, P.: DINOv2: Learning Robust Visual Features without Supervision, Transactions on Machine Learning Research, 2024.
- Ouali et al. (2020) Ouali, Y., Hudelot, C., and Tami, M.: Semi-Supervised Semantic Segmentation With Cross-Consistency Training, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 12 671–12 681, 2020.
- Rahnemoonfar et al. (2021) Rahnemoonfar, M., Chowdhury, T., Sarkar, A., Varshney, D., Yari, M., and Murphy, R. R.: FloodNet: A High Resolution Aerial Imagery Dataset for Post Flood Scene Understanding, IEEE Access, 9, 89 644–89 654, 2021.
- Rahnemoonfar et al. (2023) Rahnemoonfar, M., Chowdhury, T., and Murphy, R.: RescueNet: A High Resolution UAV Semantic Segmentation Dataset for Natural Disaster Damage Assessment, Scientific Data, 10, 913, 10.1038/s41597-023-02799-4, 2023.
- Ronneberger et al. (2015) Ronneberger, O., Fischer, P., and Brox, T.: U-Net: Convolutional Networks for Biomedical Image Segmentation, in: Medical Image Computing and Computer-Assisted Intervention – MICCAI 2015, pp. 234–241, 2015.
- Sandhini Putri et al. (2022) Sandhini Putri, A. F., Widyatmanti, W., and Umarhadi, D. A.: Sentinel-1 and Sentinel-2 data fusion to distinguish building damage level of the 2018 Lombok Earthquake, Remote Sensing Applications: Society and Environment, 26, 100 724, 2022.
- Selvaraju et al. (2017) Selvaraju, R. R., Cogswell, M., Das, A., Vedantam, R., Parikh, D., and Batra, D.: Grad-CAM: Visual Explanations from Deep Networks via Gradient-Based Localization, in: 2017 IEEE International Conference on Computer Vision (ICCV), pp. 618–626, 2017.
- Shen et al. (2022) Shen, Y., Zhu, S., Yang, T., Chen, C., Pan, D., Chen, J., Xiao, L., and Du, Q.: BDANet: Multiscale Convolutional Neural Network With Cross-Directional Attention for Building Damage Assessment From Satellite Images, IEEE Transactions on Geoscience and Remote Sensing, 60, 1–14, 2022.
- Shi et al. (2015) Shi, X., Chen, Z., Wang, H., Yeung, D.-Y., Wong, W.-k., and Woo, W.-c.: Convolutional LSTM Network: a machine learning approach for precipitation nowcasting, in: Proceedings of the 29th International Conference on Neural Information Processing Systems, vol. 1, p. 802–810, 2015.
- Stramondo et al. (2006) Stramondo, S., Bignami, C., Chini, M., Pierdicca, N., and Tertulliani, A.: Satellite radar and optical remote sensing for earthquake damage detection: results from different case studies, International Journal of Remote Sensing, 27, 4433–4447, 2006.
- Sun et al. (2021) Sun, Y., Lei, L., Guan, D., and Kuang, G.: Iterative Robust Graph for Unsupervised Change Detection of Heterogeneous Remote Sensing Images, IEEE Trans. Image Process., 30, 6277–6291, 2021.
- Sun et al. (2024a) Sun, Y., Lei, L., Guan, D., Wu, J., Kuang, G., and Liu, L.: Image Regression With Structure Cycle Consistency for Heterogeneous Change Detection, IEEE Transactions on Neural Networks and Learning Systems, 35, 1613–1627, 2024a.
- Sun et al. (2024b) Sun, Y., Wang, Y., and Eineder, M.: QuickQuakeBuildings: Post-Earthquake SAR-Optical Dataset for Quick Damaged-Building Detection, IEEE Geoscience and Remote Sensing Letters, 21, 1–5, 2024b.
- Tarvainen and Valpola (2017) Tarvainen, A. and Valpola, H.: Mean teachers are better role models: Weight-averaged consistency targets improve semi-supervised deep learning results, in: Advances in Neural Information Processing Systems, edited by Guyon, I., Luxburg, U. V., Bengio, S., Wallach, H., Fergus, R., Vishwanathan, S., and Garnett, R., vol. 30, 2017.
- Tong et al. (2012) Tong, X., Hong, Z., Liu, S., Zhang, X., Xie, H., Li, Z., Yang, S., Wang, W., and Bao, F.: Building-damage detection using pre- and post-seismic high-resolution satellite stereo imagery: A case study of the May 2008 Wenchuan earthquake, ISPRS Journal of Photogrammetry and Remote Sensing, 68, 13–27, 2012.
- Tsai et al. (2018) Tsai, Y.-H., Hung, W.-C., Schulter, S., Sohn, K., Yang, M.-H., and Chandraker, M.: Learning to Adapt Structured Output Space for Semantic Segmentation, in: 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 7472–7481, 2018.
- UNDRR (2018a) UNDRR: Economic Losses, Poverty & Disasters: 1998-2017, Report, United Nations Office for Disaster Risk Reduction (UNDRR), https://www.undrr.org/publication/economic-losses-poverty-disasters-1998-2017, 2018a.
- UNDRR (2018b) UNDRR: Words into Action guideline: Man-made/technological hazards, Report, United Nations Office for Disaster Risk Reduction (UNDRR), https://www.undrr.org/publication/words-action-guideline-man-made/technological-hazards, 2018b.
- van der Maaten and Hinton (2008) van der Maaten, L. and Hinton, G.: Visualizing Data using t-SNE, Journal of Machine Learning Research, 9, 2579–2605, 2008.
- Vu et al. (2019) Vu, T.-H., Jain, H., Bucher, M., Cord, M., and Pérez, P.: ADVENT: Adversarial Entropy Minimization for Domain Adaptation in Semantic Segmentation, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 2512–2521, 2019.
- Wang et al. (2023) Wang, D., Zhang, Q., Xu, Y., Zhang, J., Du, B., Tao, D., and Zhang, L.: Advancing Plain Vision Transformer Toward Remote Sensing Foundation Model, IEEE Transactions on Geoscience and Remote Sensing, 61, 1–15, 2023.
- Wang et al. (2024) Wang, J., Guo, H., Su, X., Zheng, L., and Yuan, Q.: PCDASNet: Position-Constrained Differential Attention Siamese Network for Building Damage Assessment, IEEE Transactions on Geoscience and Remote Sensing, 62, 1–18, 2024.
- Wang and Jin (2012) Wang, T.-L. and Jin, Y.-Q.: Postearthquake Building Damage Assessment Using Multi-Mutual Information From Pre-Event Optical Image and Postevent SAR Image, IEEE Geoscience and Remote Sensing Letters, 9, 452–456, 2012.
- Watanabe et al. (2016) Watanabe, M., Thapa, R. B., Ohsumi, T., Fujiwara, H., Yonezawa, C., Tomii, N., and Suzuki, S.: Detection of damaged urban areas using interferometric SAR coherence change with PALSAR-2, Earth, Planets and Space, 68, 131, 2016.
- Xia et al. (2023) Xia, H., Wu, J., Yao, J., Zhu, H., Gong, A., Yang, J., Hu, L., and Mo, F.: A Deep Learning Application for Building Damage Assessment Using Ultra-High-Resolution Remote Sensing Imagery in Turkey Earthquake, International Journal of Disaster Risk Science, 14, 947–962, 2023.
- Xie et al. (2021) Xie, E., Wang, W., Yu, Z., Anandkumar, A., Alvarez, J. M., and Luo, P.: SegFormer: Simple and Efficient Design for Semantic Segmentation with Transformers, in: Advances in Neural Information Processing Systems (NIPS), vol. 34, pp. 12 077–12 090, 2021.
- Xie et al. (2016) Xie, S., Duan, J., Liu, S., Dai, Q., Liu, W., Ma, Y., Guo, R., and Ma, C.: Crowdsourcing Rapid Assessment of Collapsed Buildings Early after the Earthquake Based on Aerial Remote Sensing Image: A Case Study of Yushu Earthquake, Remote Sensing, 8, 2016.
- Xue et al. (2024) Xue, Z., Zhang, X., Prevatt, D. O., Bridge, J., Xu, S., and Zhao, X.: Post-hurricane building damage assessment using street-view imagery and structured data: A multi-modal deep learning approach, arXiv preprint arXiv:2404.07399, 2024.
- Yamaguchi (2012) Yamaguchi, Y.: Disaster Monitoring by Fully Polarimetric SAR Data Acquired With ALOS-PALSAR, Proceedings of the IEEE, 100, 2851–2860, 2012.
- Yamazaki and Matsuoka (2007) Yamazaki, F. and Matsuoka, M.: Remote Sensing Technologies in Post-Disaster Damage Assessment, Journal of Earthquake and Tsunami, 01, 193–210, 2007.
- Yang et al. (2023) Yang, J., Qian, H., Xu, Y., Wang, K., and Xie, L.: Can We Evaluate Domain Adaptation Models Without Target-Domain Labels?, arXiv preprint arXiv:2305.18712, 2023.
- Yang and Soatto (2020) Yang, Y. and Soatto, S.: FDA: Fourier Domain Adaptation for Semantic Segmentation, in: IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), pp. 4084–4094, 2020.
- Ye et al. (2024) Ye, S., Zhu, Z., and Suh, J. W.: Leveraging past information and machine learning to accelerate land disturbance monitoring, Remote Sensing of Environment, 305, 114 071, 2024.
- Ye et al. (2017) Ye, Y., Shen, L., Hao, M., Wang, J., and Xu, Z.: Robust Optical-to-SAR Image Matching Based on Shape Properties, IEEE Geoscience and Remote Sensing Letters, 14, 564–568, 2017.
- Ye et al. (2019) Ye, Y., Bruzzone, L., Shan, J., Bovolo, F., and Zhu, Q.: Fast and Robust Matching for Multimodal Remote Sensing Image Registration, IEEE Transactions on Geoscience and Remote Sensing, 57, 9059–9070, 2019.
- Yonezawa and Takeuchi (2001) Yonezawa, C. and Takeuchi, S.: Decorrelation of SAR data by urban damages caused by the 1995 Hyogoken-nanbu earthquake, International Journal of Remote Sensing, 22, 1585–1600, 2001.
- Yusuf et al. (2001) Yusuf, Y., Matsuoka, M., and Yamazaki, F.: Damage assessment after 2001 Gujarat earthquake using Landsat-7 satellite images, Journal of the Indian Society of Remote Sensing, 29, 17–22, 2001.
- Zhang et al. (2023) Zhang, J., Liu, K., and Wang, M.: Flood detection using Gravity Recovery and Climate Experiment (GRACE) terrestrial water storage and extreme precipitation data, Earth System Science Data, 15, 521–540, 2023.
- Zheng et al. (2021) Zheng, Z., Zhong, Y., Wang, J., Ma, A., and Zhang, L.: Building damage assessment for rapid disaster response with a deep object-based semantic change detection framework: From natural disasters to man-made disasters, Remote Sens. Environ., 265, 112 636, 2021.