### v3

Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (Bright) to support AI-based all-weather disaster response. To the best of our knowledge, Bright is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of human-made disasters across 14 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in this dataset, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In the experiments, we have tested seven advanced AI models trained with Bright to validate the transferability and robustness. Beyond that, it also serves as a challenging benchmark for a variety of tasks under disaster scenarios, including unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM). The experimental results serve as baselines to inspire future research and model development. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as new disaster data becomes available. Bright also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven methods in support of people in disaster-affected areas.
[…]
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods have affected 4.4 billion people and caused approximately 1.3 million deaths. These disasters have also resulted in economic losses of US\$2,647 billion in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions to minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, preventing efficient on-site assessments. Earth observation (EO), also known as remote sensing, provides a safe and efficient way to obtain information on building damage in disaster areas because of its wide field of view and contactless operation.
Figure 1: An example of the wildfire occurring in Maui, Hawaii, USA, August 2023. (a) Pre-event optical imagery (© Maxar). (b) Post-event optical image (© Maxar) with land-cover features obscured by wildfire smoke. (c) Post-event SAR imagery (© Capella Space) unaffected by smoke, showing the disaster area.
The EO technologies commonly used for assessing building damage after disasters are optical and synthetic aperture radar (SAR). Optical imagery is a primary source for building damage assessment because of its intuitive and easy-to-interpret nature. For example, moderate-resolution optical EO data from the Landsat series and Sentinel-2 have been used to assess building damage (Yusuf et al., 2001; Fan et al., 2019; Sandhini Putri et al., 2022). Landsat and Sentinel-2 data are limited in spatial resolution and only provide broad approximations of affected areas, which lack accuracy for specific buildings crucial for timely rescue. The new generation of very-high-resolution (VHR) optical sensors, such as IKONOS and WorldView, provide EO data with spatial resolutions at a meter and a sub-meter level, enabling finer assessments at the level of individual buildings (Freire et al., 2014). These data have been used successfully for the assessment of building damage (Yamazaki and Matsuoka, 2007; Tong et al., 2012; Freire et al., 2014).
While accurate building damage maps can be obtained by human-expert visual interpretation of optical images, this process is time-consuming and labor-intensive for rapid large-scale assessments. In addition, it requires trained professionals. Therefore, recent literature has focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, machine learning (ML) and deep learning (DL) techniques have significantly improved the efficiency and accuracy of building damage assessment. Earlier work was toward a single disaster event with labels annotated for a particular disaster area to train a model, which is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) of different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Chen et al., 2022a; Shen et al., 2022; Kaur et al., 2023; Guo et al., 2024; Wang et al., 2024; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damages of buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1c is not affected by smoke and clearly shows the buildings damaged by the wildfire. Due to the advantages of SAR imagery, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate accurate building damage mapping compared to optical imagery (Adriano et al., 2021; Xia et al., 2025). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010; Adriano et al., 2021). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This is a significant challenge that remains and needs to be addressed.
[…]
Table 1: Comparison of Bright with the existing building damage assessment datasets. The OA indicates whether the dataset is open access (OA) or not, and GSD is an acronym for ground sampling distance (GSD). Note that since some datasets integrate other datasets, we summarize only the largest one to avoid duplication here. For example, the BDD dataset (Adriano et al., 2021) includes the Tohoku-Earthquake-2011 dataset (Bai et al., 2018) and Palu-Tsunami-2018 dataset (Adriano et al., 2019).
Dataset OA Modality GSD (m/pixel) No. of events Disaster type No. of building Granularity
ABCD (Fujita et al., 2017) Yes Optical EO 0.4 1 Tsunami N/A Image-level
(Nguyen et al., 2017) Yes Images on social media N/A 4 3 natural disasters N/A Image-level
(Cheng et al., 2021) Yes UAV image N/A 1 Hurricane 1,802 Image-level
(Xue et al., 2024) Yes Street-view image N/A 1 Hurricane 2,468 Image-level
FloodNet (Rahnemoonfar et al., 2021) Yes UAV image N/A 1 Flood 6,675 Pixel-level
RescueNet (Rahnemoonfar et al., 2023) Yes UAV image N/A 1 Hurricane 10,903 Pixel-level
Ida-BD (Kaur et al., 2023) No Optical EO 0.5 1 Hurricane 18,083 Pixel-level
xBD (Gupta et al., 2019) Yes Optical EO <0.8 15 6 natural disasters >700,000 Pixel-level
QQB (Sun et al., 2024b) Yes Optical and SAR EO <1 1 Earthquake 4,029 Pixel-level
BDD (Adriano et al., 2021) No Optical and SAR EO 1.2-3.3 9 3 natural disasters 123,453 Pixel-level
Bright Yes Optical and SAR EO 0.3-1 14 5 natural disasters 2 human-made disasters 384,596 Pixel-level
0.1 Comparison with existing datasets
The comparison between Bright and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to limited available disaster events with corresponding open-source EO data and annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. However, the xBD includes only optical EO data. It does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models. The dataset most similar to Bright is the BDD proposed by Adriano et al. (2021). The main differences between BDD and Bright datasets are: (1) Bright covers more disaster events and building instances, including both natural and human-made disasters. (2) Bright has higher spatial resolution; while the highest resolution of SAR images in BDD is 1.2 meters, Bright provides finer detail with spatial resolution from 0.3 meters to 1 meter to enable the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, Bright is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not concerned with building damage assessment.
0.2 Main contribution
The contributions of this paper are twofold:
  1. 1.
    We present Bright, the first multimodal building damage dataset with sub-meter spatial resolution, which is publicly available to the community. Bright employs a combination of pre-event optical imagery and post-event SAR imagery with various disaster events and rich geographic diversity to support the study of AI-based multimodal building damage mapping, especially in developing countries.
  2. 2.
    We evaluate a suite of contemporary models on Bright to establish robust baselines. Beyond supervised deep learning, Bright can support a wide range of AI-based methods. It enables research in UDA, SSL, UMCD, and UMIM, among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with source code and pretrained model weights, are publicly released to provide strong baselines and accelerate future developments for disaster response in the community.
1 Study areas and disaster events
Figure 2: Geographic distribution of disaster events present in the Bright.
We selected 14 disaster events across the globe for Bright, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts.
Table 2: Summary of the basic information of the Bright dataset, with disaster events listed in chronological order.
Disaster area Type of disaster Date GSD (m/pixel) Data provider / source No. of tiles No. of building
Beirut, Lebanon Explosion (EP) 04 Aug. 2020 1 Maxar &\&& Capella 133 25,496
Bata, Equatorial Guinea Explosion (EP) 07 Mar. 2021 0.5 Maxar &\&& Capella 107 8,893
Goma, DR Congo Volcano eruption (VE) 22 May 2021 0.33 Maxar &\&& Capella 123 18,741
Les Cayes, Haiti Earthquake (EQ) 14 Aug. 2021 0.48 Maxar &\&& Capella 73 18,918
La Palma, Spain Volcano Eruption (VE) 19 Sept. 2021 - 13 Dec. 2021 0.3-0.35 IGN (Spain) &\&& Capella 933 30,239
Boulder, USA Wildfire (WF) 30 Dec. 2021 - 01 Jan., 2022 0.6 NAIP &\&& Capella 77 8,365
Ukraine Armed conflict (AC) 22 Mar. 2022 - 21 Sept. 2022 0.6 Google Earth &\&& Capella 513 56,770
Turkey Earthquake (EQ) 06 Feb. 2023 0.30-0.35 Maxar &\&& Capella &\&& Umbra 1,114 135,033
Kyaukpyu, Myanmar Cyclone (CC) 14 May 2023 0.6 Google Earth &\&& Capella 126 8,052
Maui, Hawaii, USA Wildfire (WF) 08 Aug. 2023 - 09 Aug. 2023 0.6 NOAA &\&& Capella 65 3,995
Morocco Earthquake (EQ) 08 Sept. 2023 0.35-0.4 Maxar &\&& Capella 567 6,269
Derna, Libya Flood (FL) 10 Sept. 2023 0.35 Maxar &\&& Capella 124 1,0979
Acapulco, Mexico Hurricane (HC) 25 Oct. 2023 0.35-0.8 Google Earth &\&& Capella 212 18,437
Noto, Japan Earthquake (EQ) 01 Jan. 2024 0.5 GSI &\&& Umbra 79 8,153
Total - - 0.3-1 - 4,246 384,596
1.1 Explosion in Beirut, 2020
On August 4, 2020, a massive explosion occurred at the Port of Beirut in Lebanon, caused by the improper storage of 2,750 tons of ammonium nitrate. The explosion caused widespread damage within a radius of several kilometers, significantly impacting the port and the surrounding neighborhoods, including areas such as Gemmayzeh, Mar Mikhael, and Achrafieh. It resulted in 218 deaths, more than 7,000 injuries, and left approximately 300,000 people homeless (Fakih and Majzoub, 2021). Economic losses were estimated to be 15 billion USD (Valsamos et al., 2021). The disaster compounded Lebanon’s ongoing economic challenges and contributed to political instability and social unrest.
1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city (OCHA, 2021a). A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
1.3 Volcano Eruption in DR Congo and Rwanda, 2021
On May 22, 2021, Mount Nyiragongo in the Democratic Republic of the Congo erupted, causing widespread devastation. The eruption resulted in 32 deaths, the destruction of 1,000 homes, and the displacement of thousands as lava flows threatened the city of Goma (IFRC, 2021). Nearly 400,000 people were evacuated due to the risk of further volcanic activity, including potential magma flow beneath Goma and nearby Lake Kivu. Despite continued seismic activity, life in Goma largely returned to normal by August 2021, although plans to relocate parts of the city remain under consideration due to the ongoing threat from the volcano.
1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed (OCHA, 2021b). As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
1.5 Volcano Eruption in La Palma, 2021
On September 19, 2021, the Cumbre Vieja volcano on La Palma, part of Spain’s Canary Islands, erupted following several days of seismic activity. The eruption primarily impacted the island’s western side, covering over 1,000 hectares with lava and destroying more than 3,000 buildings, including the towns of Todoque and La Laguna. The lava flow, measuring about 3.5 kilometers wide and 6.2 kilometers long, reached the sea, cutting off the coastal highway and forming a new peninsula with extensive lava tubes. Although the timely evacuation of around 8,000 people prevented major casualties, one person died from inhaling toxic gases (Troll et al., 2024). Economic losses exceeded 800 million Euros, and thousands of residents were displaced, with significant damage to arable land and livelihoods.
1.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the state history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over US\$2 billion in damage (Forrister et al., 2024). More than 37,500 residents evacuated, and significant damage to public drinking water systems occurred (Forrister et al., 2024).
1.7 Armed Conflict in Ukraine, 2022
In February 2022, Russian forces launched a full-scale invasion of Ukraine, resulting in widespread destruction and displacement. By November 2024, total damages to Ukraine’s infrastructure had reached \$170 billion. More than 236,000 residential buildings were damaged or destroyed, including 209,000 private houses, 27,000 apartment buildings, and 600 dormitories. Over 4,000 educational institutions and 1,554 medical facilities were also affected, alongside extensive damage to transport, energy, and telecommunications infrastructure. The conflict has inflicted more than 40,000 civilian casualties, displaced four million people internally, and forced 6.8 million refugees to flee. As of late 2024, approximately 14.6 million Ukrainians require humanitarian assistance (Andrienko et al., 2025).
1.8 Earthquake in Turkey, 2023
On February 6, 2023, a magnitude 7.8 earthquake struck southeastern Turkey near Gaziantep, followed by a magnitude 7.7 aftershock. The disaster, the most powerful earthquake in Turkey since 1939, caused widespread destruction across approximately 350,000 $km^{2}$, affecting 14 million people and displacing 1.5 million. The death toll reached 53,537, with 107,213 injuries (STL, 2024), making it one of the deadliest earthquakes in modern history. Economic losses were estimated at USD 148.8 billion, with over 518,000 houses and 345,000 apartments destroyed (Government of Türkiye, 2023). The earthquake caused severe damage to infrastructure, agriculture, and essential services, further worsening the region’s economic challenges. International aid was mobilized to support the affected populations.
1.9 Cyclone in Myanmar, 2023
In May 2023, Cyclone Mocha, a Category 5 hurricane, struck Myanmar, causing widespread devastation in the country’s coastal regions, particularly in Rakhine State. According to official estimates, at least 148 people were killed and 132 injured, although other sources suggest higher figures (GFDRR and World Bank, 2023). The cyclone affected around 1.2 million people in Rakhine alone, with over 200,000 buildings reportedly damaged or destroyed, making it one of the most destructive cyclones in the region in the past 15 years. The direct economic damage was estimated at US\$2.24 billion in damages, equivalent to 3.4% of Myanmar’s GDP (GFDRR and World Bank, 2023).
1.10 Wildfire in Hawaii, 2023
In August 2023, a series of wildfires broke out on the island of Maui, Hawaii, causing widespread destruction and significant impacts on the local population and environment. The fires, fueled by dry conditions and strong winds, primarily affected the town of Lahaina, where at least 102 people were killed and two remain missing (Hedayati et al., 2024). Over 2,200 buildings were destroyed, including many historic landmarks, resulting in estimated damages of US\$5.5 billion (Jones et al., 2024; NCEI, 2025). The fires prompted evacuations and led to the displacement of thousands of residents, with significant economic losses in the tourism and agriculture sectors.
1.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The earthquake killed nearly 3,000 people, injured more than 5,500 and displaced more than half a million (OCHA, 2023). The earthquake damaged or destroyed nearly 60,000 houses, with the heaviest losses reported in rural communities of the Atlas Mountains (International Federation of Red Cross and Red Crescent Societies, 2024). Overall damage is estimated at about USD 7 billion (National Geophysical Data Center / World Data Service, 2025), while direct economic losses of roughly USD 30 million amount to approximately $0.24\%$ of Morocco’s GDP. Cultural heritage in Marrakesh was also hard-hit: sections of the UNESCO-listed Medina and several historic mosques sustained severe damage.
1.12 Flood and Storm in Libya, 2023
In September 2023, Storm Daniel brought catastrophic flooding to northeastern Libya, particularly in Derna, when two dams collapsed and released an estimated 30 million cubic meters of water. At least 5,923 people were killed, though local officials warned fatalities could reach 18,000–20,000 (USAID Bureau for Humanitarian Assistance, 2024). Post–event analyses revealed that approximately $10\%$ of Derna’s housing stock was destroyed and a further $18.5\%$ sustained damage. Across the wider coastal belt—from Benghazi, through Jabal Al Akhdar and Al Marj, to Derna—an estimated 18 838 dwellings were damaged or obliterated (Normand and Heggy, 2024).. The disaster, considered the second deadliest dam failure in history, devastated the city’s infrastructure: four bridges collapsed, and entire neighborhoods were washed out to sea. Long-standing neglect of dam maintenance, compounded by Libya’s political turmoil, contributed significantly to the scale of the tragedy.
1.13 Hurricane in Mexico, 2023
In September 2023, Hurricane Norma, a Category 4 hurricane, struck the western coast of Mexico, severely affecting Sinaloa and Baja California Sur. This was followed by Hurricane Otis in October, which made landfall near Acapulco as a Category 5 hurricane. Otis was the strongest Pacific hurricane to hit Mexico, causing at least 52 deaths and leaving 32 missing. The storm caused unprecedented destruction, with more than 51,864 homes destroyed and damages estimated at US\$12–16 billion, surpassing Hurricane Wilma as the costliest Mexican hurricane (Reinhart and Reinhart, 2024).
1.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 7.45-meter tsunami along the Sea of Japan coast. The disaster resulted in 572 deaths and over 1,300 injuries, with 193,529 structures damaged across nine prefectures (Fire and Disaster Management Agency, 2025). It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20, 12,929 people remained in 521 evacuation centers (Japanese Red Cross Society, 2024). The event prompted Japan’s first major tsunami warning since 2011. Total damage is estimated 16.9 trillion yen, reflecting substantial economic losses with significant impacts on infrastructure and communities across the affected regions (Kaneko, 2024).
Figure 3: Overall flowchart of developing the Bright dataset.
Figure 4: Thumbnails of local areas in 14 disaster events of Bright. For visualization purposes, different events have different scales.
2 Dataset Description
2.1 Construction of the dataset
Figure 3 shows the flowchart of developing Bright. The optical EO data in our dataset are mainly from Maxar’s Open Data program [^3] https://www.maxar.com/open-data, while the SAR EO data are from Capella Space [^4] https://www.capellaspace.com/earth-observation/gallery and Umbra [^5] https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consists of red, blue, and green bands, while the SAR EO data are amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the data to an 8-bit data format. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two types of EO data. Figure 11 shows the selected control points on three disaster scenes.
Table 3: Definition of the different categories of damage in Bright.
Category Definition
Background (0) All non-building pixels
Intact (1) No visible signs of structural damage, water intrusion, shingle displacement, or burn marks.
Damaged (2) Partial structural damage to the building, such as missing roof members, visible cracks, or partial wall/roof collapse. Buildings may be partially burned, surrounded by water or mud, or affected by nearby volcanic flows.
Destroyed (3) Completely collapsed, burned, partially/completely covered by water/mud or no longer present
The labels in Bright consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections of EO experts to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service [^6] https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products [^7] https://unosat.org/products, and The Federal Emergency Management Agency (FEMA) [^8] https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To further improve damage information reliability, we performed a secondary validation and correction of these annotations. EO experts carefully reviewed the damage points using multi-temporal imagery available on platforms such as Google Earth and manually refined the labels when inconsistencies or misclassifications were observed. Since the number of damage levels varied across different events, we refered to the framework in FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019) to standardize them into three categories: intact (1), damaged (2), and destroyed (3). Table 3 summarizes these definitions. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to possible coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
Figure 4 presents thumbnails of selected local areas from the 14 disaster events.
2.2 Statistics for Bright
Figure 5: Statistics for the Bright dataset. (a) Distributions of band values of samples from four study sites. (b) Distribution of building scales. (c) Feature distribution of buildings on four events under two imaging modalities. (d) Percentage of building and background pixels and percentage of different damage levels in building pixels.
The basic information about Bright, including disaster events, EO data, the number of corresponding EO tiles, and the total number of building pixels, is summarized in Table 2. After cropping the EO data into 1024$\times$1024-pixel tiles, Bright contains 4,246 multimodal image pairs.
The key statistics of Bright are illustrated in Figure 5. Figure 5a shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting Bright’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5b shows that Bright covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
[…]
Table 4: Proxy registration errors (in pixel) estimated using different multimodal image registration methods. The estimated registration errors for each event are reported in Table 13.
Descriptor RMSE (All) RMSE (Building)
RIFT (Li et al., 2020) 1.125 1.138
LNIFT (Li et al., 2022) 0.857 0.825
SRIF (Li et al., 2023a) 1.090 1.054
Average 1.024 1.006
We also evaluated the registration accuracy between optical and SAR EO data in Bright. Due to the absence of real ground truth, we introduce a proxy method to estimate registration errors, leveraging existing multimodal image descriptors. This approach is detailed in Appendix B. Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements, obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within building regions. Accurate registration results ensure spatial consistency across modalities, providing a solid foundation for accurate building damage assessment.
2.3 Dataset division strategy
To train DL models using Bright and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) and checkpoints are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
In addition to the above standard machine learning division, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events, a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. This setup comprises two key divisions:
  • Zero-shot setup: This setting mimics a real-world scenario where a newly occurring disaster must be analyzed without any prior labeled data from the same event. We isolate one event as an unseen test set while using the remaining 13 events for training and validation. This setting evaluates the cross-event generalization ability of models, testing how well they can transfer learned knowledge from previous disasters to an entirely new event. Due to the high variability of disaster types and geographies, this setup is inherently challenging, as models trained on past disasters may struggle to assess damage patterns in a previously unseen event accurately.
  • One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup, acknowledging that in real-world disaster response, a small number of labeled samples from the new disaster can be obtained. In this setting, a limited subset of labeled data (one pair for training and one pair for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.
3 Methodology
3.1 Problem statement and two solutions
The objective of building damage assessment is to interpret EO data covering the affected area and generate a building damage map that reflects the extent of damage to buildings. To achieve this, two common solutions are typically employed.
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted. This process can be formalized as $\mathbf{Y}^{dam}=\mathcal{M}^{seg}(\mathbf{X}^{\mathrm{T}_{1}},\mathbf{X}^{\mathrm{T}_{2}})$, where is $\mathbf{X}^{\mathrm{T}_{1}}$ is the pre-event imagery, $\mathbf{X}^{\mathrm{T}_{2}}$ is the post-event imagery, $\mathcal{M}^{seg}\left(\cdot\right)$ is a semantic segmentation model, $\mathbf{Y}^{dam}$ is the obtained damage map. In the context of this paper, $\mathbf{X}^{\mathrm{T}_{1}}$ is VHR optical imagery and $\mathbf{X}^{\mathrm{T}_{2}}$ is VHR SAR imagery.
The second solution adopts the task decoupling approach (Gupta et al., 2019; Zheng et al., 2021), which breaks down building damage assessment into two subtasks: the building localization task, i.e., separating the building from the background, and the damage classification task, i.e., focusing on the classification between different levels of damage. This solution can be formulated as $\mathbf{Y}^{loc}=\mathcal{M}^{loc}(\mathbf{X}^{\mathrm{T}_{1}})$ and $\mathbf{Y}^{clf}=\mathcal{M}^{clf}(\mathbf{X}^{\mathrm{T}_{1}},\mathbf{X}^{\mathrm{T}_{2}})$, where $\mathbf{Y}^{loc}$ is the building localization map, $\mathbf{Y}^{clf}$ is the damage classification map, $\mathcal{M}^{loc}\left(\cdot\right)$ and $\mathcal{M}^{clf}\left(\cdot\right)$ are models for building localization and damage classification tasks, respectively. $\mathcal{M}^{loc}\left(\cdot\right)$ and $\mathcal{M}^{clf}\left(\cdot\right)$ can be two separate models (Gupta et al., 2019) or a unified multi-task learning model (Zheng et al., 2021; Chen et al., 2022a, 2024). The final building damage map is obtained by combining the two outputs using a simple mask operation: $\mathbf{Y}^{dam}=\mathbf{Y}^{loc}\odot\mathbf{Y}^{clf}$.
Since the aim of this article is not only to provide a large-scale multimodal dataset to support all-weather disaster response but also to offer insights for designing appropriate methods in future research, we will employ and compare both solutions in our experiments.
3.2 Benchmark suites
We evaluate several advanced deep network architectures from both the computer vision and EO communities on Bright. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of representative DL architectures, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), Transformers, and the more recent Mamba architecture.
Among these seven networks, UNet, DeepLabV3+, and SiamAttnUNet adopt the first solution, directly treating building damage assessment as a single semantic segmentation task. In contrast, SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba adopt the second solution by decoupling the task into building localization and damage classification tasks.
Beyond supervised deep learning, Bright also enables the evaluation of broader strategies commonly explored in EO and vision communities:
  • Unsupervised domain adaptation (UDA) methods for the zero-shot transfer setup, enabling models to transfer knowledge across disaster events with no labeled samples from the target event.
  • Semi-supervised learning (SSL) approaches for the one-shot transfer setup, leveraging a small number of labeled samples and the remaining unlabeled samples from new disaster events to refine model adaptation.
  • Unsupervised multimodal change detection (UMCD) methods, which exploit the modality-independent relationship in optical and SAR data to detect land-cover changes without requiring manual annotations.
  • Unsupervised multimodal image matching (UMIM) methods, which aim to learn modality-independent features to enable automatic registration of multimodal data without relying on manual alignment.
Table 5: Accuracy assessment for different DL models on the test set under the standard machine learning division (set-level mIoU). The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
[…]
Background Intact Damaged Destroyed
UNet 87.97 72.24 95.47 64.94 96.19 71.27 39.13 53.17
DeepLabV3+ 87.00 70.33 95.43 64.80 95.98 70.98 37.53 54.69
SiamAttnUNet 88.16 70.13 95.45 64.26 96.14 71.90 35.06 53.92
SiamCRNN 89.45 72.02 95.76 65.73 96.48 73.44 38.91 54.18
ChangeOS 89.60 71.88 95.84 65.98 96.54 73.85 38.99 54.53
DamageFormer 90.29 72.51 96.13 67.09 96.87 75.04 39.86 56.59
ChangeMamba 90.90 72.70 96.22 67.63 96.96 75.59 40.05 57.91
3.3 Model training
To train these supervised models, we use a combination of cross-entropy loss and Lovasz softmax loss (Berman et al., 2018). Cross-entropy loss serves as the basic loss function for dense prediction tasks, while Lovasz softmax loss effectively addresses sample imbalance between non-building and building pixels and across different damage levels.
For UNet, DeepLabV3+, and SiamAttnUNet, which directly predict damage maps from the input multimodal image pairs, the training loss function is defined as:
$\mathcal{L}^{bda}_{coupled}=\mathcal{L}^{bda}_{ce}+\mathcal{L}^{bda}_{lov}.$ (1)
For SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba, which decouple building damage assessment into building localization and damage classification subtasks, the training loss function is defined as:
$\mathcal{L}^{bda}_{decoupled}=\mathcal{L}^{loc}_{ce}+\mathcal{L}^{loc}_{lov}+\mathcal{L}^{clf}_{ce}+\mathcal{L}^{clf}_{lov}.$ (2)
[…]
For zero-shot and one-shot setups, we incorporate UDA and SSL techniques to better utilize unlabeled and limited target disaster data. The specifics of these training adjustments, along with the UMCD and UMIM approaches, are provided in Appendix E, F, G, and H, respectively.
Table 6: The mIoU on different events for different DL models (event-level mIoU). The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
Events UNet DeepLabV3+ SiamAttnUNet SiamCRNN ChangeOS DamageFormer ChangeMamba
Beirut-EP-2020 52.23 48.53 51.85 52.04 52.57 60.55 49.38
Bata-EP-2021 39.24 40.26 38.69 41.44 39.80 40.37 40.47
Goma-VE-2021 61.52 59.26 62.95 62.50 61.84 61.55 62.89
Les Cayes-EQ-2021 40.39 38.52 40.84 40.99 41.24 41.75 42.36
La Palma-VE-2021 62.72 64.31 65.56 64.24 65.34 64.27 65.13
Marshall-WF-2021 53.96 44.67 51.48 56.57 56.55 57.50 57.23
Ukraine-AC-2022 43.92 43.84 40.71 45.52 45.56 47.45 47.43
Turkey-EQ-2023 49.34 49.53 48.57 50.69 52.34 51.36 52.70
Kyaukpyu-CC-2023 38.11 38.56 40.49 42.35 42.03 48.42 41.05
Hawaii-WF-2023 53.02 57.81 57.92 55.34 58.40 60.82 60.70
Morocco-EQ-2023 42.62 42.05 42.58 43.17 43.05 44.00 44.06
Derna-FL-2023 59.47 56.08 66.82 59.48 57.73 62.24 63.30
Acapulco-HC-2023 49.21 46.66 42.15 49.98 51.08 50.82 47.01
Noto-EQ-2024 36.49 37.33 36.56 38.35 42.84 40.57 45.78
Average 48.73 47.67 49.08 50.19 50.74 52.26 51.39
3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), the F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the building damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
4 Results and Analysis
4.1 Evaluation results on standard machine learning split
[…]
Figure 6: Damage maps predicted by different models on the test images of 14 events. The meaning of the color in reference maps and damage maps is consistent with Figure 4.
Finally, Figure 6 shows some of the building damage maps obtained by the seven models on the test set.
4.2 What have models learned and what can they learn?
Figure 7: Visualization of feature response to “Damaged” and “Destroyed” categories in different layers of deep models over three event cases. (a) Goma-VE-2021. (b) Beirut-EP-2023. (c) Libya-FL-2023. (d) Hawii-WF-2023. (e) Turkey-EQ-2023. (f) Acapulco-HC-2023. In the visualization, closer to red indicates larger response values, and closer to dark blue the opposite. See Appendix A for implementation details.
To better understand model behavior beyond performance metrics, we explore the internal attention patterns of trained models using class activation maps (CAMs). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events, including volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. Taking Goma-VE-2021 as the example, for the “Destroyed” category, the encoder processing the optical imagery exhibits strong activations across nearly all built-up regions, accurately localizing individual buildings. This suggests the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, the encoder handling the SAR imagery shows intense activation over the lava-covered region on the left, indicating that the model has identified this as a key signal for destruction—likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, however, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder stage, the destroyed buildings in the lava-affected area are strongly activated, aligning well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as the subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
Figure 8: IoU distribution of deep models over seven disaster types.
[…]
In summary, these findings confirm both the promise and limitations of Optical–SAR combinations for all-weather, global-scale disaster response. While the combination performs well in events characterized by large-scale surface disruption (e.g., wildfires, volcanoes), it struggles with subtle or localized damage patterns. Incorporating richer data sources, such as fully polarimetric SAR and LiDAR, can further enhance the accuracy and reliability of future building damage assessments.
Table 7: Further contributions to mIoU from post-processing algorithms. ChangeMamba (Chen et al., 2024) is used here as the baseline. Details on these algorithms are provided in Appendix D.
Method mIoU (set) mIoU (event)
Baseline 67.63 51.39
[…]
4.3 Effect of post-processing methods
Post-processing techniques help refine the raw predictions from DL models, reducing noise, improving consistency, and ensuring spatial coherence in damage maps (Zheng et al., 2021). Here, we further explore the impact of post-processing algorithms. Table 7 presents the effect of three post-processing techniques applied to ChangeMamba. The post-processing methods evaluated include test-time augmentation, object-based major voting, and model ensembling. Further details on these methods are provided in Appendix D.
[…]
In summary, post-processing techniques play a crucial role in improving damage proxy maps. Future work can further explore adaptive post-processing strategies tailored to specific disaster types to enhance prediction reliability in multimodal EO data contexts.
Table 8: The mIoU on different events for different DL models in zero-shot and one-shot cross-event transfer setup. The highest values are highlighted in purple, and the second-highest results are highlighted in teal.
Events UNet DeepLabV3+ SiamAttnUNet SiamCRNN ChangeOS DamageFormer ChangeMamba
zero- / one-shot zero- / one-shot zero- / one-shot zero- / one-shot zero- / one-shot zero- / one-shot zero- / one-shot
Beirut-EP-2020 31.62 / 34.89 34.72 / 33.58 35.33 / 29.17 39.43 / 40.94 38.48 / 38.21 38.71 / 39.93 40.30 / 40.92
Bata-EP-2021 33.48 / 37.48 33.55 / 30.23 33.78 / 35.24 37.46 / 40.06 38.00 / 38.18 38.91 / 39.42 39.17 / 40.31
Goma-VE-2021 55.97 / 57.14 55.01 / 58.13 57.52 / 55.75 58.91 / 59.54 58.66 / 57.85 55.86 / 59.89 60.39 / 61.91
Les Cayes-EQ-2021 26.21 / 38.72 35.50 / 37.90 37.22 / 38.76 39.45 / 40.28 39.36 / 40.26 40.41 / 41.07 41.35 / 42.06
La Palma-VE-2021 32.74 / 35.29 30.86 / 36.25 34.00 / 34.99 32.90 / 33.32 29.96 / 32.60 32.28 / 37.62 34.18 / 34.93
Marshall-WF-2021 34.97 / 37.84 35.05 / 37.12 35.23 / 39.54 39.82 / 46.83 36.31 / 43.72 40.93 / 51.15 40.61 / 45.96
Ukraine-AC-2022 32.36 / 37.50 32.23 / 35.98 33.84 / 35.76 34.77 / 38.14 36.30 / 39.28 33.84 / 36.51 35.78 / 36.99
Turkey-EQ-2023 37.82 / 35.83 37.28 / 40.49 36.09 / 37.49 38.32 / 40.11 38.45 / 39.51 37.20 / 40.27 41.28 / 42.80
Kyaukpyu-CC-2023 31.46 / 31.71 34.01 / 34.75 33.77 / 35.93 34.19 / 37.71 36.66 / 36.25 32.99 / 33.54 33.85 / 36.08
Hawaii-WF-2023 31.73 / 39.80 31.32 / 45.44 31.79 / 45.12 30.94 / 42.65 31.15 / 41.96 31.60 / 50.64 36.14 / 50.89
Morocco-EQ-2023 38.24 / 39.81 40.48 / 41.23 40.78 / 40.69 40.89 / 42.89 41.40 / 41.89 41.89 / 43.42 42.86 / 43.51
Derna-FL-2023 33.25 / 38.44 35.07 / 40.02 33.76 / 42.15 35.88 / 44.77 35.80 / 43.85 36.27 / 44.69 37.02 / 45.47
Acapulco-HC-2023 26.99 / 38.76 27.78 / 33.24 30.93 / 38.94 26.70 / 34.53 27.06 / 34.50 27.48 / 32.87 28.45 / 38.96
Noto-EQ-2024 31.02 / 33.83 35.57 / 39.17 35.87 / 36.22 36.04 / 38.52 39.14 / 39.15 36.37 / 43.93 39.90 / 44.33
Average 34.13 / 38.36 35.60 / 38.82 36.42 / 38.98 37.55 / 41.45 37.62 / 40.52 37.48 / 42.49 39.38 / 43.23
4.4 Evaluation results on cross-event transfer setup
4.4.1 Baseline methods
[…]
4.4.2 Unsupervised domain adaptation and semi-supervised learning methods
Table 9: Evaluation results of some unsupervised domain adaptation methods on zero-shot cross-event transfer setup. DeepLabV3+ is used as the baseline here. Detailed mIoU on each event is listed in Table 14.
Method Avg. mIoU
Source-Only 35.60
[…]
Table 10: Evaluation results of some semi-supervised learning methods on one-shot cross-event transfer setup. DeepLabV3+ is used as the baseline here. Detailed mIoU on each event is listed in Table 15.
Method Avg. mIoU
Baseline 38.82
[…]
In the context of cross-event transfer, UDA and SSL methods naturally emerge as promising strategies to bridge the domain gap between source and target disaster events. In the zero-shot setting, UDA methods aim to improve model generalization by aligning the source and target domains without requiring any target labels. In the one-shot setting, SSL methods further leverage a small number of labeled samples along with abundant unlabeled data from the target event, making them especially appealing for real-world disaster scenarios where rapid annotation is infeasible. However, it remains an open question whether these UDA and SSL methods, originally developed for natural image domains, can effectively handle the challenges of multimodal EO data in complex disaster scenarios. To examine the applicability of these methods, we evaluate several representative UDA and SSL algorithms using DeepLabV3+ as the baseline model. The results are reported in Tables 9 and 10, respectively.
[…]
4.4.3 Why is cross-event transfer challenging?
Figure 9: The comparison of pixel distribution of different categories in SAR images for different events. The first row compares different events with the same disaster type. (a) Explosion. (b) Volcano eruption. (c) Wildfire. (d) Cyclone/Hurricane (e) Earthquake. The second row compares different disaster types. (f) Explosion vs Conflict. (g) Wildfire vs Volcano Eruption. (h) Cyclone vs Flood. (i) Explosion vs Wildfire. (j) Conflict vs Earthquake.
To better understand why cross-event generalization is so difficult, we explore two fundamental factors rooted in the nature of real-world disaster response.
  1. 1.
    Inconsistent damage signatures across events. Figure 9 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. Firstly, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are not consistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. Secondly, the distribution shift becomes more pronounced across different disaster types (the second row). For instance, the signatures of destroyed building in wildfires are notably different from those in volcano eruptions, which again differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster may learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
    Figure 10: Comparison of models’ best performance (mIoU) on test events versus the best checkpoints selected on validation sets under the cross-event transfer setting. The farther a point lies from the diagonal line, the larger the gap between the model’s selected performance and its true upper bound.
  2. 2.
    Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a target domain validation set is available to tune hyperparameters or select the best checkpoint. However, in disaster response scenarios, no such validation data can be assumed. Figure 10 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future solutions should consider both aspects, i.e., developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).
4.5 Extended evaluation of unsupervised multimodal change detection methods
Table 11: Evaluation results of representative unsupervised multimodal change detection methods. The highest values are highlighted in purple, and the second-highest results are highlighted in teal. The accuracies on the UMCD benchmark dataset are the accuracies on the four datasets presented in Figure 12. Details of methods and benchmark datasets are introduced in Appendix G.
Method UMCD benchmark datasets Bright
OA F1 KC OA F1 IoU KC
IRG-McS (Sun et al., 2021) 98.3 / - / 97.1 / 97.2 80.4 / - / 75.4 / 73.7 79.4 / - / 73.9 / 75.1 90.03 12.65 6.75 7.65
SR-GCAE (Chen et al., 2022b) 98.6 / 98.5 / - / - 82.9 / 77.6 / - / - 82.1 / 76.9 / - / - 77.83 14.35 7.73 5.64
FD-MCD (Chen et al., 2023) 98.2 / 97.8 / - / 96.7 81.4 / 72.2 / - / 73.2 82.3 / 71.1 / - / 71.4 80.96 15.84 8.60 7.94
AOSG (Han et al., 2024) - / - / - / 96.4 - / - / - / 77.7 - / - / - / 75.9 77.93 10.75 5.68 3.98
AGSCC (Sun et al., 2024a) 98.3 / - / 95.9 / 97.6 78.2 / - / 68.0 / 77.9 77.3 / - / 65.8 / 76.6 88.49 14.82 8.00 9.54
AEKAN (Liu et al., 2025) 98.7 / - / - / 98.3 83.8 / - / - / 84.7 83.1/ - / - / 83.9 81.60 13.09 7.00 3.56
Unsupervised multimodal change detection (UMCD) plays a crucial role in post-disaster assessment by enabling rapid analysis of affected areas without requiring labeled data or prior model training (Lv et al., 2022). The current research is mainly conducted on toy datasets with limited geographic diversity and scene complexity. These datasets mainly involve general land-cover changes. It is still unknown whether these methods can work in real disaster occurrence/building damage scenarios. Bright offers a new opportunity to evaluate UMCD methods, thereby providing insights into their robustness and scalability. Six representative UMCD methods (Sun et al., 2021; Chen et al., 2022b, 2023; Han et al., 2024; Sun et al., 2024a; Liu et al., 2025) are evaluated in this paper. The specific experimental setup is described in Appendix G.
[…]
  1. 1.
    Limitations in traditional UMCD datasets. The traditional UMCD datasets consist of only a handful of image pairs, often depicting simple land-cover changes such as urban expansion or deforestation, with relatively low-resolution imagery with limited geographic diversity, as shown in Figure 12. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, Bright provides thousands of VHR multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
  2. 2.
    Interference from non-disaster changes. Unlike previous UMCD benchmarks where general land-cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, Bright requires methods to differentiate true building structural damage from irrelevant land-cover variations, posing a unique challenge to current approaches.
  3. 3.
    Problematic evaluation protocol. Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real disaster scenario, there are no labeled samples available for hyperparameter tuning, making the current practice unrealistic. Bright exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results highlight the challenges of applying existing UMCD models to real-world disaster scenarios. Bright reveals significant limitations in current methodologies and also presents new opportunities for future UMCD research.
4.6 Extended evaluation of unsupervised multimodal image matching methods
Table 12: Registration performance of different UMIM methods on the scene of Noto-Earthquake-2024. Note that the offsets here are in meters and are calculated based on the geo-coordinates of control points manually selected by EO experts. N/A means that the method is unable to complete the registration task.
Method Before (meters) After (meters)
FLSS (Ye et al., 2017) 16.97 10.01
HOPC (Ye et al., 2019) 9.37
LNIFT (Li et al., 2022) N/A
SRIF (Li et al., 2023a) N/A
Finally, precise image alignment is a critical prerequisite for any multimodal EO application. Bright offers a unique opportunity to evaluate the performance of existing UMIM algorithms under realistic, large-scale disaster conditions. Due to the lack of real pixel-level ground truth correspondences, we adopt a proxy evaluation strategy using manually selected control points (as shown in Figure 11) as references. These points are selected by several EO experts to represent clearly identifiable and stable features across modalities. While this does not constitute absolute ground truth, the use of such human-verified correspondences provides a valuable reference, allowing us to assess how closely automated methods approximate human matching ability under multimodal and disaster conditions.
Table 12 presents the quantitative results of four UMIM methods on the Noto-Earthquake-2024 scene. These methods fall into two main categories: feature-based methods (LNIFT (Li et al., 2022) and SRIF (Li et al., 2023a)), which rely on sparse keypoint detection and matching, and area-based methods (FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019)), which operate on local regions to estimate correspondences. A detailed description of each method is provided in Appendix H. Our experimental results reveal that traditional feature-based methods generally fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable. In contrast, area-based methods, which rely on the similarity in a local region rather than point correspondence in the global level, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that the Bright benchmark and its challenging real-world settings will inspire the development of new UMIM methods that are more robust to large scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
5 Discussion
5.1 Limitation of Bright
The performance of ML/DL models heavily depends on the quality of the training data. The Bright dataset is designed to enable ML/DL models trained on it to generate building damage maps of practical significance for future disaster events. It is anticipated that Bright will serve as a benchmark for numerous subsequent studies and practical disaster relief applications. However, the Bright dataset also has some limitations.
  1. 1.
    Registration error. Bright involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 4.
  2. 2.
    Label quality. The building polygons in Bright were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on Bright. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise.
  3. 3.
    Sample and regional imbalance. Although Bright is richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 (1,114 tiles) v.s. Hawaii-Wildfire-2023 (65 tiles) in Table 2. Although we used an additional event-level evaluation method, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. It may affect the generalizability of the trained model in real disaster scenarios. Apart from that, all disaster events are located near the equator or in the northern hemisphere, with no events from the southern hemisphere represented. This spatial bias could potentially limit the applicability of trained models to regions with different building styles, land cover patterns, or SAR imaging characteristics prevalent in the southern part of the globe.
Overall, despite the limitations, it is the first time such a multimodal VHR dataset has been constructed that caters to multimodal EO research under large-scale and diverse disaster context.
5.2 Significance of Bright
Timely disaster response is often hindered by delays in both EO data acquisition and damage interpretation workflows (Ye et al., 2024). Traditional expert-driven building damage mapping is time-consuming and not scalable. While ML and DL methods offer automated alternatives, their effectiveness remains limited by the scope and quality of available training data. Most existing open-source datasets are optical-only, restricting the models’ operational applicability in adverse weather and low-light conditions. As the first globally distributed multimodal dataset, Bright encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on Bright to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, Bright offers several distinct characteristics: multimodal data, VHR imagery with sub-meter spatial resolution, coverage of five types of natural disasters, as well as two kinds of human-made disasters, rich geographic diversity, and open access to the community. Due to these features, Bright is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications.
Beyond building damage assessment, Bright can also be used to support several research directions within the EO and vision community. In this work, we have applied Bright to evaluate the performance of several UDA and SSL methods for cross-event transfer. Additionally, we have demonstrated its applicability to UMCD and UMIM, showcasing its versatility as a benchmark for multiple EO challenges under real-world constraints. Furthermore, the dataset provides a strong foundation for broader multimodal EO research. Its high-quality annotations and geographic diversity make it directly applicable to tasks such as building footprint extraction, land cover mapping, height estimation, and even EO-based visual question answering (VQA). Researchers can also repurpose or extend Bright to create task-specific benchmarks, enabling flexible experimentation across tasks. Bright is also well-positioned to support the development of remote sensing foundation models, large-scale, pre-trained models designed to generalize across sensors, tasks, and regions (Wang et al., 2023; Hong et al., 2024). Its rich combination of modalities, spatial detail, and contextual diversity provides the kind of data diversity required to build such general-purpose, transferable models. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024)
Looking ahead, Bright can be further enhanced by incorporating additional modalities, such as elevation data or multi-band SAR/optical inputs, to support terrain-aware analysis and improve semantic reasoning. Future versions may also help fill current geographic gaps, including southern hemisphere regions, ensuring more globally representative coverage. Ultimately, We envision that Bright, true to its name, will bring even a glimmer of brightness to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
[…]
In this paper, we introduced Bright, the first globally distributed, multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. Bright includes pre-event optical imagery and post-event SAR imagery with sub-meter spatial resolution. Beyond introducing the dataset, we conducted a comprehensive series of experiments to validate its utility. We benchmarked several state-of-the-art supervised learning models under a standard machine learning split, and further extended the evaluation to a cross-event transfer setting, simulating real-world scenarios where no/limited target annotations are available. Additionally, we assessed the performance of unsupervised domain adaptation and semi-supervised learning methods, as well as unsupervised multimodal change detection and image matching techniques. These findings provide both performance baselines and valuable insights into model design for future research. Bright is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve Bright’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.
[…]
The Bright dataset and the code for training and testing benchmark methods (including code related to IEEE GRSS DFC 2025) are available at https://github.com/ChenHongruixuan/BRIGHT.
Appendix A Feature visualization
For the feature visualisation of buildings in different events in Figure 5c, we specifically employ DINOv2 (Oquab et al., 2024) to extract the high-dimensional features of buildings in different events. The corresponding reference maps are used to exclude background features. The features are then visualized using the t-SNE algorithm (van der Maaten and Hinton, 2008).
To visualize the spatial focus of models during building damage mapping, we apply Grad-CAM technique [^9] https://github.com/jacobgil/pytorch-grad-cam, a gradient-based technique for generating class-specific activation maps (Selvaraju et al., 2017). We perform Grad-CAM analysis on the ChangeOS model. Specifically, we compute the activation responses for the “Damaged” and “Destroyed” categories from the following components: the third or fourth ResNet block of the two branches of encoder, capturing mid- to high-level modality-specific features; the fusion module of the damage classification branch in the decoder, reflecting early-stage fusion and decision-making. These activations help reveal which spatial regions the model relies on when assessing different damage levels and how it utilizes information from optical and SAR modalities differently, depending on the disaster context.
Appendix B Manual registration and estimating registration errors
Figure 11: Overview of manually selected control points used for registration on some events. The number in parentheses of each event is the number of control points selected on that view of the image. The labeled yellow boxes highlight specific zoom-in regions that are shown in below.
The manually selected control points by EO experts on some disaster scenes are shown in Figure 11. Despite multiple rounds of meticulous registration and cross-validation by several EO experts, registration errors between optical and SAR images cannot be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we propose a proxy method for registration error estimation:
  1. 1.
    Feature point selection: Select many feature points (e.g., 3000 points) from both the optical and SAR tiles using a keypoint detection algorithm.
  2. 2.
    Feature extraction: Use validated and well-performing multimodal registration descriptors (Li et al., 2020, 2022, 2023a) to extract modality-independent features for the selected points.
  3. 3.
    Feature matching: Perform feature matching between the extracted feature points across the optical and SAR images.
  4. 4.
    Registration error computation: After completing the matching process, compute the pixel distances between the registered point pairs. These distances serve as a proxy for registration error.
  5. 5.
    Outlier removal: To improve the reliability of the estimation, we exclude matched points with a root mean square error (RMSE) greater than a certain threshold (e.g., 20 pixels). This threshold is set because after multiple rounds of expert corrections and cross-checking, large registration errors have already been eliminated. Points exceeding this threshold can be considered outliers.
Table 13: The estimated proxy registration error (in pixel) for each event.
Event RMSE (All) RMSE (Building)
Beirut-EP-2020 1.365 1.365
Bata-EP-2021 1.226 1.160
Goma-VE-2021 0.966 0.975
Les Cayes-EQ-2021 1.392 1.377
La Palma-VE-2021 0.976 0.944
Marshall-WF-2021 1.263 1.266
Ukraine-AC-2022 0.899 0.882
Turkey-EQ-2023 0.937 0.901
Kyaukpyu-CC-2023 0.970 0.998
Hawaii-WF-2023 1.124 1.124
Morocco-EQ-2023 1.088 1.050
Derna-FL-2023 1.267 1.255
Acapulco-HC-2023 1.166 1.148
Noto-EQ-2024 1.072 1.043
By following this method, we can estimate the registration errors between optical and SAR images in the dataset, providing a reliable foundation for further damage assessment analysis.
Appendix C Details of benchmark DL models
Here, we present some implementation details of our baseline model.
  • For UNet, the encoder has five convolutional blocks, each consisting of two 3$\times$3 convolutional layers and two batch normalization layers. The number of channels in the five blocks are 64, 128, 256, 512, and 1024, respectively. The decoder contains four convolutional blocks, and the number of channels is set accordingly to the number of channels of the skip-connected features extracted by the encoder.
  • For DeepLabV3+, we applied ResNet-50 initialized with ImageNet pretraining weights (He et al., 2016) as the encoder. We then modified the input channel of the encoder so that the network can predict building damage maps directly from the optical and SAR stacked images.
  • For SiamAttnUNet, we followed the settings in its original literature (Adriano et al., 2021).
  • For SiamCRNN, we applied pseudo-siamese ResNet-18 as encoder and four ConvLSTM layers (Shi et al., 2015) with 3$\times$3 convolutional kernel size and a hidden dimension of 128, as the decoder.
  • For ChangeOS, we used the ResNet-18 initialized with ImageNet pretraining weights as the encoder. We modified the encoder to a pseudo-siamese structure to extract features from pre-event and post-event images with different modalities.
  • For DamageFormer, we used pseudo-siamese Swin-Transformer-Tiny (Liu et al., 2021) as the encoder instead of pure-siamese MixFormer (Xie et al., 2021) as in the original literature (Chen et al., 2022a).
  • For the ChangeMamba family of models (Chen et al., 2024), we chose one of them, MambaBDA-Tiny, for our experiments. We modified the encoder to a pseudo-siamese structure to extract features from pre-event and post-event images with different modalities.
Appendix D Details of post-processing methods
To further refine the raw predictions from deep learning models and enhance the quality of building damage maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 4.3. This section provides an overview of these methods and details the specific implementation used in our study.
  • Test-time augmentation is a widely used technique to improve model robustness by applying transformations to the input images at inference time. Instead of making a single prediction per image, multiple augmented versions of the same input are passed through the network, and the resulting predictions are aggregated. This reduces model sensitivity to spatial variations and increases prediction stability. In our study, we apply rotation (90°, 180°, 270°) and horizontal/vertical flipping to the input images during inference. The network produces a set of predictions for each augmented version, which are then aggregated at the logit level by summing the outputs before applying the softmax function.
  • Object-based major voting aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
  • Model ensembling is a technique that combines predictions from multiple models to reduce uncertainty and improve generalization. By leveraging diverse models with different architectures, ensembling helps smooth out individual model biases and enhances the overall robustness of predictions. We perform ensembling by combining the results of three top-performing models: ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). Each model independently generates a damage proxy map, and the final prediction is obtained by averaging their logits before applying the softmax function.
Appendix E Details of unsupervised domain adaptation methods
Table 14: The mIoU on different events for different UDA methods in Table 9.
Events AdaptSeg AdvEnt CLAN PyCDA FDA
Beirut-EP-2020 35.55 33.07 37.25 32.15 35.89
Bata-EP-2021 37.54 35.20 36.98 30.01 36.31
Goma-VE-2021 52.49 51.54 56.91 49.67 52.46
Les Cayes-EQ-2021 37.11 36.86 39.47 33.40 33.90
La Palma-VE-2021 30.54 30.04 30.10 32.76 30.50
Marshall-WF-2021 37.20 35.01 34.35 30.14 33.93
Ukraine-AC-2022 32.36 32.65 32.21 32.92 35.03
Turkey-EQ-2023 33.69 34.54 36.68 34.58 36.82
Kyaukpyu-CC-2023 34.62 29.22 30.83 31.31 36.71
Hawaii-WF-2023 36.52 42.17 33.36 34.30 36.19
Morocco-EQ-2023 37.93 37.42 39.04 30.55 38.76
Derna-FL-2023 36.69 35.31 35.24 37.31 38.53
Acapulco-HC-2023 27.05 27.32 27.73 30.11 31.68
Noto-EQ-2024 35.49 35.65 33.99 25.44 31.39
Average 36.05 35.43 36.01 33.19 36.29
To evaluate the effectiveness of UDA techniques in the cross-event zero-shot setting, we selected several representative methods from the computer vision literature originally designed for semantic segmentation tasks. The following methods are implemented:
  • AdaptSeg (Tsai et al., 2018): an adversarial training framework that aligns output space distributions between source and target domains.
  • AdvEnt (Vu et al., 2019): a variant of adversarial adaptation focusing on entropy minimization and structured output alignment.
  • CLAN (Luo et al., 2019): a category-level alignment method that selectively aligns class-specific features.
  • PyCDA (Lian et al., 2019): a curriculum domain adaptation framework that iteratively refines pseudo-labels across domains.
  • FDA (Yang and Soatto, 2020): a simple style-transfer-based approach that transfers source images into the target domain appearance via frequency mixing.
All methods are implemented using DeepLabV3+ as the backbone architecture. This choice is motivated by the fact that DeepLab-based networks are commonly used in the official implementations of these UDA methods, and many codebases are tightly coupled with DeepLab’s structure. Therefore, adopting DeepLabV3+ as the shared backbone ensures compatibility with existing implementations and avoids the need for extensive code modification. It also provides a fair and consistent basis for comparison across all methods.
In our experimental setup, each test disaster event is treated as the target domain, and only its unlabeled optical-SAR image pairs are used for adaptation. Importantly, no labeled target domain data is used for training or model selection, in line with our zero-shot assumption. This setup is fundamentally different from conventional UDA protocols in the computer vision community, where a target domain validation set is typically available to select the best model checkpoint, and the target test set is held out purely for evaluation. In contrast, our setting is designed to reflect real-world disaster response scenarios, where it is realistic to obtain unlabeled imagery, but no ground-truth annotations are available at adaptation time. Therefore, model selection must rely solely on source domain feedback, which introduces additional challenges and more closely aligns with the operational constraints of emergency response.
The hyperparameters for each method closely follow the original papers to preserve method fidelity. Below are additional implementation details specific to certain methods:
  • FDA: For each training iteration, we generate frequency-transferred source images using FDA and mix them 1:1 with the original source images. The combined dataset is used to train the model from scratch.
  • PyCDA: Training proceeds in two stages. First, the model is trained on the source domain for 30,000 iterations. In the second stage, the adaptation module is applied to the target domain using pseudo-labeling and uncertainty-aware refinement as proposed in the original work.
For all methods, the same data preprocessing pipeline and training schedule (e.g., learning rate, batch size) are used unless otherwise specified in the original method, like those supervised models. The final evaluation is conducted on the target domain, and performance is reported in terms of mIoU.
Appendix F Details of semi-supervised learning methods
Table 15: The mIoU on different events for different semi-supervised learning methods in Table 10.
Events MT CCT GCT CPS
Beirut-EP-2020 32.91 35.68 35.48 35.46
Bata-EP-2021 37.94 36.61 38.25 36.96
Goma-VE-2021 56.46 57.68 53.52 47.49
Les Cayes-EQ-2021 40.20 39.57 38.83 39.41
La Palma-VE-2021 33.52 33.03 36.90 33.29
Marshall-WF-2021 39.73 40.12 38.74 32.94
Ukraine-AC-2022 39.11 36.34 36.44 31.34
Turkey-EQ-2023 38.96 39.72 36.50 37.15
Kyaukpyu-CC-2023 36.72 35.84 33.56 30.22
Hawaii-WF-2023 54.10 48.93 49.99 47.64
Morocco-EQ-2023 37.18 40.54 41.10 37.54
Derna-FL-2023 37.55 41.07 40.45 40.28
Acapulco-HC-2023 32.30 30.16 33.36 31.51
Noto-EQ-2024 43.40 37.61 37.74 37.65
Average 40.00 39.49 39.34 37.06
In the one-shot cross-event transfer setting, we evaluate several representative SSL methods for semantic segmentation. These methods were originally developed for natural image domains and are adapted here to the multimodal EO-based building damage assessment scenario.
  • Mean Teacher (MT) (Tarvainen and Valpola, 2017) maintains an exponential moving average (EMA) of model weights to generate stable predictions for unlabeled data, enforcing consistency between the student and teacher outputs.
  • CCT (Ouali et al., 2020) applies perturbations to the input data and enforces consistency between different views of the same image using multiple decoder branches.
  • GCT (Ke et al., 2020) further enhances this idea by introducing guidance from a pre-trained model to regularize learning.
  • CPS (Chen et al., 2021) trains two separate networks that generate pseudo-labels for each other, encouraging cross-supervised learning on unlabeled data.
All methods also use DeepLabV3+ as the backbone network. This decision is consistent with their original implementations and allows for seamless integration with official codebases without the need for architectural modifications. It also ensures a fair and consistent comparison across methods.
In the setup, 13 labeled disaster events serve as the training and validation sets, while a single labeled sample from the test (target) event is provided to simulate a realistic one-shot adaptation scenario. In addition, the remaining unlabeled samples from the test event are made available and used to facilitate SSL.
Each model is trained in two stages. In the first stage, the model is trained for 10,000 iterations using only standard supervised losses, including cross-entropy and Lovász-Softmax on the labeled data from the source domain and the single labeled sample from the target domain. This stage serves to initialize the model with a stable representation before semi-supervised training begins. In the second stage, each SSL method applies its respective semi-supervised objective using the unlabeled data from the target domain. These include consistency regularization (e.g., Mean Teacher), cross-view perturbation training (e.g., CCT and GCT), or dual-network pseudo supervision (e.g., CPS). Hyperparameters for each method follow the configurations suggested in their original publications.
Final performance is evaluated on the test set of the target domain using mIoU as the primary evaluation metric.
Appendix G Details of unsupervised multimodal change detection methods and a more practical evaluation protocol
Figure 12: Illustration of benchmark datasets commonly used in unsupervised multimodal change detection. The order from left to right is the corresponding order in the Table 11. In reference maps, white color indicates changed pixels, and black color indicates unchanged pixels.
We further evaluate several recent UMCD methods on Bright. These methods are designed to detect binary changes from paired multimodal imagery without relying on manual labels.
  • IRG-McS (Sun et al., 2021) constructs superpixel-based graphs for each modality and iteratively matches their structural relationships to detect changes across heterogeneous image pairs, followed by Markov co-segmentation for final change map generation.
  • SR-GCAE (Chen et al., 2022b) learns robust graph-based representations of local and nonlocal structural relationships in multimodal images using a graph convolutional autoencoder, enabling effective change detection without supervision.
  • FD-MCD (Chen et al., 2023) proposes a Fourier domain framework that analyzes both local and nonlocal structural relationships using graph spectral convolution and adaptive fusion, enabling robust unsupervised multimodal change detection.
  • AOSG (Han et al., 2024) constructs an adaptively optimized structured graph to capture patch-level structural features in multimodal images, iteratively refining change intensity measures by fusing self-change and cross-domain structural differences for accurate unsupervised change detection.
  • AGSCC (Sun et al., 2024a) translates multimodal images via a structure cycle-consistent image regression framework that enforces similarity in structural graphs across domains, using adaptive graphs and multiple regularization terms to robustly detect changes without supervision.
  • AEKAN (Liu et al., 2025) utilizes a superpixel-based Siamese AutoEncoder built on Kolmogorov–Arnold Networks (KAN) to extract latent commonality features between multimodal images, using reconstruction and hierarchical consistency losses to detect changes in an unsupervised manner.
The training and evaluation protocols commonly adopted in the UMCD literature are often limited to individual image pairs (shown in Figure 12), using the same scene for representation learning, hyperparameter tuning, and evaluation. While this is understandable due to the lack of large-scale public datasets in this field, such a setup fails to reflect real-world use cases and often leads to overfitting and overestimated performance. To address this, we introduce a standardized and practical evaluation protocol using the Bright dataset. Specifically, we use the validation set from the standard ML split in Section 2.3 as the training set for UMCD methods, including any hyperparameter or threshold tuning. These models are then evaluated on the combined training and test sets, which are strictly held out during adaptation. This avoids overlap between tuning and evaluation scenes and better simulates the real-world deployment setting, where a model is expected to generalize to new, unseen disasters.
For evaluation, it is important to note that most of these methods were originally developed for land cover change detection, where changes correspond to transitions between semantic categories such as vegetation, water, or built-up areas. Applying them directly to disaster damage detection poses certain challenges. In particular, “Damaged” buildings do not always exhibit strong spatial or spectral signals, making them hard to distinguish in an unsupervised setting. However, “Destroyed” buildings often result in a complete change of land cover appearance (e.g., collapsed structures, debris, or ground exposure), which aligns better with the assumptions of existing UMCD methods. Therefore, in our evaluation, we treat buildings labeled as “Destroyed” in Bright as the positive (changed) class, and all other regions (including intact, damaged, and background) as the unchanged class. This allows a fair adaptation of these methods to a disaster-relevant setting while respecting their original design.
Appendix H Details of unsupervised multimodal image matching methods
In our extended evaluation of UMIM methods, we categorized the selected algorithms into two groups based on their matching strategies: feature-based and area-based methods.
Feature-based methods, such as LNIFT (Li et al., 2022) and SRIF (Li et al., 2023a), aim to detect sparse keypoints independently in both optical and SAR images and compute modality-invariant feature descriptors. The matching is then performed globally by comparing descriptor similarity between the two modalities. These methods are implemented using publicly available code [^10] https://github.com/LJY-RS/SRIF.
In contrast, area-based methods, including FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019), begin by identifying keypoints only in the optical image. These points are then matched by sliding a template window across a local search region in the SAR image, relying on local patch similarity. In our experiments, we used a template size of 120$\times$120 pixels, and a search window of 200$\times$200 pixels for both methods. This category is also implemented using open-source code [^11] https://github.com/yeyuanxin110/CFOG.
We performed registration on the entire original EO image pair from the Noto-Earthquake-2024 event. Because of the large spatial size of the input images, images were downsampled to half their original resolution to improve computational efficiency and maintain feasibility for all methods.
For quantitative evaluation, we adopted a control-point-based proxy metric. A set of manually selected control points, validated by EO experts shown in Figure 11, is used to represent stable, cross-modal, and clearly identifiable features (e.g., building corners, road intersections). The average spatial offset is computed using their coordinates in a common projected coordinate system (e.g., UTM), both before and after registration. Before registration, the offset can be calculated as:
$E_{\text{before}}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(x_{i}^{\text{opt}}-x_{i}^{\text{sar}})^{2}+(y_{i}^{\text{opt}}-y_{i}^{\text{sar}})^{2}}.$ (3)
where $(x_{i}^{\text{opt}},y_{i}^{\text{opt}})$ and $(x_{i}^{\text{sar}},y_{i}^{\text{sar}})$ are the projected coordinates of the $i$-th control point in the optical and SAR images, respectively.
After registration, SAR points are transformed using the estimated mapping $\mathcal{T}$, and the post-registration error is computed as:
$E_{\text{after}}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(x_{i}^{\text{opt}}-\hat{x}_{i}^{\text{sar}})^{2}+(y_{i}^{\text{opt}}-\hat{y}_{i}^{\text{sar}})^{2}},$ (4)
where $(x_{i}^{\text{sar}},y_{i}^{\text{sar}})=\mathcal{T}(x_{i}^{\text{opt}},y_{i}^{\text{opt}})$ are transformed SAR coordinates.
This approach provides an alternative approximation of alignment accuracy in the absence of ground truth. Although this metric is not a substitute for full correspondence maps, it offers a practical and interpretable measure of registration performance, and enables us to assess how well unsupervised methods can approach human-level matching under complex, real-world disaster conditions.
[…]
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (equal). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (lead). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (equal). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan [^12] Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).
[…]
The contact author has declared that none of the authors has any competing interests.
Acknowledgements.
This work was supported in part by the JSPS, KAKENHI under Grant Number 24KJ0652 and 22H03609, the Council for Science, Technology and Innovation (CSTI), the Cross-ministerial Strategic Innovation Promotion Program (SIP), Development of a Resilient Smart Network System against Natural Disasters (Funding agency: NIED), JST, FOREST under Grant Number JPMJFR206S, Microsoft Research Asia, Next Generation AI Research Center of The University of Tokyo, and Young Researchers Exchange Programme between Japan and Switzerland under the Japanese-Swiss Science and Technology Programme. The authors would also like to give special thanks to Sarah Preston of Capella Space, Capella Space’s Open Data Gallery, Maxar Open Data Program, and Umbra’s Open Data Program for providing the valuable data.